### v1

# Allowed Coulomb branch scaling dimensions of four-dimensional $\mathcal{N}=2$ SCFTs

Abstract
A basic datum of a rank-$r$ $\mathcal{N}{=}2$ superconformal field theory (SCFT) is the $r$-tuple of its Coulomb branch scaling dimensions, i.e., the scaling dimensions of a set of special protected scalar operators whose vevs generate the coordinate ring of the Coulomb branch of the theory. It is well known that when the coordinate ring is freely generated these scaling dimensions can only take values in a small set of rational numbers. But there are further constraints on which $r$-tuples of these numbers can appear. The main aim of this work is to clarify what these are. Along the way we also compute explicitly the $r$-tuples of allowed scaling dimensions for theories of ranks $r=2,3,4$.
1 Introduction and summary
The space of unitary four-dimensional $\mathcal{N}{=}2$ superconformal field theories (SCFTs) of non-zero rank appears to be remarkably constrained by the possible Coulomb branch (CB) geometries of their moduli spaces, characterized and constrained by their rigid special Kähler (SK) structures and scaling symmetries. When a rank-$r$ (i.e., $r$-complex dimensional) CB is free of complex singularities, its coordinate ring is freely generated by the vacuum expectation values (vevs) $\{u_{i},\ i=1,\dots,r\}$ of a special set of operators, the Coulomb branch operators, with definite scaling dimensions $\{\Delta_{i}\}$. The problem of determining which $r$-tuples of CB dimensions are allowed is the subject of this paper.
It has been shown that the scaling dimensions for a rank-$r$ CB are restricted to take values in the finite set of rational numbers Caorsi:2018zsq ; Argyres:2018urp
[…]
where $\varphi$ is the Euler totient function. (We re-derive this result in section 3.) However, not all sets of $r$-many such values give rise to consistent CB geometries. In fact, the allowed $r$-tuples of dimensions represent a significantly smaller subset, thus making the CB scaling dimension $r$-tuple an even handier tool to preliminarily determine the consistency of a given candidate rank $r$ $\mathcal{N}{=}2$ SCFT, heavily restricting the possible theories in question. For example, setting $r=2$ in (1) gives the 23 allowed dimensions
$\Delta_{j}^{r=2}\in\left\{\frac{12}{11},\frac{10}{9},\frac{8}{7},\frac{6}{5},\frac{5}{4},\frac{4}{3},\frac{10}{7},\frac{3}{2},\frac{8}{5},\frac{5}{3},\frac{12}{7},2,\frac{12}{5},\frac{5}{2},\frac{8}{3},3,\frac{10}{3},4,5,6,8,10,12\right\}.$ (2)
Thus the naïve expectation is that there should be 276 distinct allowed pairs while a closer analysis cuts down this number to 63.111This number differs from the 79 pairs stated in Kaidi:2022sng . This is because that reference did not account for the non-reflexivity of the genuine scaling dimension conditions explained here in section 3.2 below equation (31). See, e.g., Cecotti:2023ksl for further examples. As such, mapping out the consistent set of scaling dimensions at a given rank provides invaluable information for the ongoing classification of $\mathcal{N}{=}2$ SCFTs.
We present here a careful and systematic construction of the full set of algebraic relations which CB operator scaling dimensions have to satisfy, from which the allowed $r$-tuples follow. The algebraic relations are a consequence of the EM duality monodromies and the structure of singular loci encoded in the special geometry of scale-invariant CBs. In particular, this clarifies previous works Caorsi:2018zsq ; Kaidi:2022sng ; Argyres:2022lah by some of the authors which contained partial or incorrect characterizations of the allowed $r$-tuples.
It is worth reminding the reader that the correspondence between $r$-tuples of CB scaling dimensions and $\mathcal{N}=2$ SCFTs is neither injective nor surjective. In fact, there are many known cases where several inequivalent $\mathcal{N}=2$ SCFTs share the same $r$-tuple of scaling dimensions (see, for example, Martone:2021ixp ). Similarly, there is no guarantee that all allowed $r$-tuples are realized. While there are $\mathcal{N}=2$ SCFTs realizing all allowed rank-1 scaling dimensions Argyres:2015ffa ; Argyres:2015gha ; Argyres:2016xmc , already at rank-2 there are several couples $(\Delta_{1},\Delta_{2})$ which have no known realisation (so far) as scaling dimensions of CB operators of a known SCFT.
This paper is organized as follows. In section 2, we review the most relevant aspects of scale-invariant CB geometries with no complex singularities, focusing on their scaling and special Kahler stratifications. This review is by no means self-contained, but is meant to bring together and highlight results from Argyres:2020wmq ; Cecotti:2023ksl . Next, in section 3, we delve into the way the two stratifications interact to constrain the allowed $r$-tuples. The resulting constraints on allowed $r$-tuples are summarized around equation (31). Finally, section 4 presents an algorithm for calculating all physically admitted sets of CB scaling dimensions. Appendix A tabulates the allowed scaling dimensions at rank-2, rank-3, and rank-4 which follow from this algorithm. For each rank we only tabulate the genuinely new tuples of scaling dimensions Martone:2021ixp ; Argyres:2022lah ; Argyres:2022puv ; Argyres:2022fwy . This is a subset of allowed scaling dimensions but it is the essential datum that allows to compute the entire set. The precise definition of this subset can be found below in section 3.1.
2 A review of Coulomb branch geometry
Other reviews of CB geometry can be found in Alvarez-Gaume:1996ohl ; Lerche:1996xu ; Freed:1997dp ; Martone:2020hvy .
2.1 Scale invariant Coulomb branches without complex singularities
Any rank-$r$ $4$d $\mathcal{N}{=}2$ SCFT possesses a space of gauge inequivelant vacua known as its moduli space.222 The moduli space has several branches, some of which special, that can be characterized by the pattern of breaking of the $\mathrm{U}(2)$ superconformal $R$-symmetry. In particular, the Higgs branch is characterized by vevs that are preserving the $\mathrm{U}(1)_{R}$ symmetry while spontaneously breaking the $\mathrm{SU}(2)_{R}$ symmetry, and the Coulomb branch is characterized by preserving the $\mathrm{SU}(2)_{R}$ symmetry and spontaneously breaking the $\mathrm{U}(1)_{R}$ symmetry. The CB of rank-$r$, denoted $\mathcal{C}$, is the branch of the moduli space in which the low energy dynamics at each point are described by a $\mathrm{U}(1)^{r}$ gauge theory. It is an $r$-complex-dimensional space with a rigid special Kähler (SK) geometry with metric non-analyticities on a subspace $\mathcal{V}\subset\mathcal{C}$ Freed:1997dp ; Seiberg:1994aj ; Seiberg:1994rs . Physically, the vacua where the CB metric is non-analytic are vacua for which the theory has extra massless charged states.
If we assume that the CB has no complex structure singularities — an assumption we will be making throughout — it is always possible to find a set of $r$ complex coordinates, $u:=(u_{1},\ldots,u_{r})$, which are globally defined on $\mathcal{C}$ and which we will take as coordinates on this space. These can also be identified with the vevs of the CB operators — scalar superconformal primaries which are singlets of $\mathrm{SU}(2)_{R}$ and which satisfy the shortening condition that their $\mathrm{U}(1)_{R}$ charge is proportional to their scaling dimension — which generate the CB chiral ring Beem:2014zpa . The absence of complex singularities is translated algebraically by the requirement that the CB chiral ring is freely generated Argyres:2017tmj . If it is not the case, we have relations between the vevs of the generating set of CB operators and a unique set of CB scaling dimensions becomes harder to define. Herein we will only consider theories with a freely generated chiral ring.
The CB inherits additional structure from the parent SCFT. The scale symmetry from the conformal symmetry group and the $\mathrm{U}(1)_{R}$ symmetry of the superconformal group act non-trivially on the CB. We define the action of dilatations on CB operators as $\Phi_{i}(x)\mapsto\lambda^{\Delta_{i}}\Phi_{i}(\lambda^{-1}x)$, for $\lambda$ a positive real number. The weights, $\Delta_{j}$, of operators under this action are called their scaling dimensions, and are constrained to be greater than or equal to 1 by unitarity. The scaling dimension is proportional to the $\mathrm{U}(1)_{R}$ charge for CB operators, so they combine to give a complex action on the CB. In particular, under the assumption of freely generated CB chiral ring this action is generated by the holomorphic vector field
$\mathcal{E}=\sum_{j}\Delta_{j}u_{j}\frac{\partial}{\partial u_{j}}$ (3)
where $u_{j}\doteq\langle\Phi_{j}\rangle$ are the vevs, assumed to be global complex coordinates on the CB. This exponentiates to a $\mathbb{C}^{\times}$ action on $\mathcal{C}$ given by
×$\mathbb{C}^{\times}:u$ $\mapsto\exp(z\mathcal{E})\cdot u\doteq\left(e^{z\Delta_{1}}u_{1},\ldots,e^{z\Delta_{r}}u_{r}\right),$ (4)
for all $z\in\mathbb{C}$.
The $r$-tuple of CB scaling dimensions, $(\Delta_{1},\ldots,\Delta_{r})$, is the main object of study of this paper. Although the $(u_{i})$ coordinates are not uniquely defined by (4) — linear redefinitions of $u_{i}$’s with the same dimension are possible — the $r$-tuple of scaling dimensions (including multiplicities) is uniquely defined.
Associated to this $\mathbb{C}^{\times}$ action is a stratification of the CB by its $(u_{i})$ coordinate hyperplanes. In particular, for each subset $(i_{1},\ldots,i_{\ell})\subset(1,\ldots,r)$, define a $-\ell$ hyperplane in $\mathcal{C}\simeq\mathbb{C}^{r}$ by
$\mathcal{I}_{i_{1},\ldots,i_{\ell}}$$\doteq\{u\in\mathbb{C}^{r}\,|\,u_{j}=0,\ \forall j\notin(i_{1},\ldots,i_{\ell})\}.$ (5)
We will call the subset of scaling dimensions $(\Delta_{i_{1}},\ldots\Delta_{i_{\ell}})$ those associated to $\mathcal{I}_{i_{1},\ldots,i_{\ell}}$. We will be particularly interested in the 1-dimensional $u_{i}$-coordinate axis $\mathcal{I}_{i}$ with associated dimension $\Delta_{i}$.
The interplay of this $\mathbb{C}^{\times}$ stratification with the stratification of the CB generated by its metric non-analyticities will be key to our investigation. This latter stratification — the SK stratification — is quite constrained Argyres:2020wmq ; Cecotti:2023ksl by the SK structure of the CB and the physical interpretation of the CB non-analyticities, as we now review. We start by reviewing the SK structure of the CB away from its metric non-analyticities.
2.2 SK geometry of the CB and its associated algebraically integrable system
The states in the theory at a generic point of the CB are specified by their charges under the $\mathrm{U}(1)^{r}$ gauge symmetry. We will collectively call these $\mathbf{p}$. Dirac quantisation restricts $\mathbf{p}$ to lie in a lattice $\Lambda\cong\mathbb{Z}^{2r}$ equipped with an antisymmetric integer pairing called the Dirac pairing, $J$. Each point on the charge lattice has a pair of electric and magnetic charges for each $\mathrm{U}(1)$ gauge factor. The group that preserves $\Lambda$ and a given Dirac pairing $J$, the electric-magnetic (EM) duality group, is $\mathrm{Sp}_{J}(2r,\mathbb{Z})$. If the pairing is principal, this is the usual symplectic group $\mathrm{Sp}(2r,\mathbb{Z})$. [^3] Theories with non-principal pairings are so-called relative theories Freed:2012bs ; Aharony:2013hda . See, e.g., Gaiotto:2010be ; DelZotto:2022ras ; Argyres:2022kon ; Closset:2023pmc for more on 4d relative QFTs. $\mathrm{Sp}_{J}(2r,\mathbb{Z})$ duality transformations leave the physics of the $\mathrm{U}(1)^{r}$ theory invariant.
Vevs of the scalar superpartners of the $\mathrm{U}(1)^{r}$ field strengths (in an “electric” EM duality frame) are special coordinates, $a^{i}(u)$, on the CB. In a magnetic duality frame, they are dual special coordinates, $a_{i}^{D}(u)$. These transform under the $\mathrm{Sp}_{J}(2r,\mathbb{Z})$ EM duality group as holomorphic vector-valued functions on the CB,
$\sigma(u)\doteq\left(\begin{smallmatrix}\mathbf{a}^{D}(u)\\ \mathbf{a}(u)\end{smallmatrix}\right).$ (6)
Unbroken $\mathcal{N}{=}2$ supersymmetry implies that the matrix of low-energy $\mathrm{U}(1)^{r}$ complex gauge couplings is given in terms of the special coordinates by
$\tau_{ij}(u)=\frac{\partial a_{i}^{D}}{\partial a^{j}}.$ (7)
The Kähler metric components on the CB (i.e., the kinetic terms of the scalars) with respect to the $a^{j}$ special coordinates are given by $g_{ij}=\text{Im}(\tau_{ij})$. This identification together with unitarity imply that the coupling matrix $\tau_{ij}$ is symmetric and has positive definite imaginary part. These structures and properties define the special Kähler (SK) structure of the CB.
The SK structure can be succinctly described in terms of the special section $\sigma$ of an $\mathrm{Sp}_{J}(2r,\mathbb{Z})$ vector bundle over $\mathcal{C}$ whose fiber, $V^{*}$, is the linear dual of the complexification of the charge lattice. The complexification of the charge lattice inherits both the Dirac pairing and the action of the EM duality group. Then the CB Kähler potential is $K=iJ(\sigma,\bar{\sigma})$, with (7) and the symmetry of $\tau_{ij}$ is ensured by the condition $J(\partial_{i}\sigma,\partial_{j}\sigma)=0$. Positive-definiteness of the Kähler metric ensures positivity of $\text{Im}(\tau_{ij})$. Note that, since the special coordinates are vevs of free scalar fields which have mass dimension 1, the weight of $\sigma(u)$ under the $\mathbb{C}^{\times}$ action is $\Delta_{\sigma}=1$. Note also that $\sigma$ does not diverge anywhere in $\mathcal{C}$, as this would give a sub-sector of the theory that is decoupled at all scales.
The seminal works of Seiberg:1994aj ; Seiberg:1994rs ; Donagi:1995am ; Donagi:1995cf associate an algebraically integrable system to the non-singular locus of the CB; that is, a fibration of polarized abelian varieties $X_{u}$ over the CB $\mathcal{C}$ endowed with a holomorphic symplectic two-form $\omega$ that vanishes when restricted to the fibers. The latter property can be summarized as saying that the fibration is Lagrangian with respect to $\omega$. Let us briefly discuss this notion, before using it to constrain the possible sets of CB scaling dimensions in section 3. For more detailed accounts of these notions, we refer the reader to Donagi:1995am ; lange2013complex and to appendix A of Argyres:2024hdn .
An abelian variety is a complex torus $A=\mathbb{C}^{r}/\Lambda$ that is also a projective variety. As such, in order to fully specify an abelian variety, we must also state how to embed the complex torus $A$ into $\mathbb{P}^{n}$ for some $n\in\mathbb{N}$. This can be achieved by equipping $A$ with an integral non-degenerate skew pairing $J$ on $\Lambda$ [^4] Geometrically, $J$ corresponds to the first Chern class of an ample line bundle $\mathcal{L}$ on $A$. As $\mathcal{L}$ is ample, the sections of an appropriate power of it defines a closed immersion of $A$ into $\mathbb{P}^{n}$ for some $n$. Such a $J$ is called a polarization on $A$ and is said to be principal if $\det J=1$. In the context of $\mathcal{N}{=}2$ SQFTs, the electromagnetic charge lattice $\Lambda_{u}$, IR effective couplings $\tau_{ij}(u)$, and the Dirac pairing $J$ provide the datum of a polarized abelian variety associated to a point $u\in\mathcal{C}$.
Concretely, define $X_{u}=\mathbb{C}^{r}/\Lambda_{u}$, and let $\{\eta_{i}\}$ be a basis of $H^{(1,0)}(X_{u})\cong\mathbb{C}^{r}$ and $\{\alpha^{i},\beta_{j}\}$ be a basis of $H_{1}(X_{u})\cong\Lambda_{u}$. Define the period matrix, $\Pi$, of $X_{u}$ with respect to these bases as the $(r\times 2r)$-dimensional matrix
$\Pi=\big{(}{\textstyle\int_{\boldsymbol{\alpha}}\boldsymbol{\eta},\int_{\boldsymbol{\beta}}\boldsymbol{\eta}}\big{)}.$ (8)
By the Riemann conditions, $X_{u}$ is an abelian variety if and only if there are bases such that $\Pi=\big{(}J,\tau\big{)}$, with $J$ and $\tau_{ij}$ satisfying the SK conditions described above. In this way the SK structure defines a holomorphic fibration of abelian varieties over the CB. Note that the non-degeneracy of $J$ and $\text{Im}\,\tau_{ij}$ imply that
$\det\left(\begin{smallmatrix}\Pi\\ \bar{\Pi}\end{smallmatrix}\right)\neq 0,$ (9)
which is just the statement that $\Lambda_{u}$ is a full rank lattice in $\mathbb{C}^{r}$, so $X_{u}$ is a non-degenerate torus.
Furthermore, if $\omega$ is a holomorphic two-form on the total space of this fibration of abelian varieties over $\mathcal{C}$ with vanishing restriction to the fibers, then its fiber integrals are well-defined. If we identify these fiber integrals with the CB differentials of the special section,
$\mathop{}\!\mathrm{d}a^{i}(u)=\int_{\alpha^{i}}\omega,\quad\mathop{}\!\mathrm{d}a^{D}_{i}(u)=\int_{\beta_{i}}\omega,$ (10)
then the SK conditions on the special coordinates follow from $\omega$ being closed and non-degenerate on the total space. This means that $\omega$ is a holomorphic symplectic form with respect to which the fibration is lagrangian. In this way the algebraically integrable system is defined by the CB SK geometry.
We view $\omega^{-1}$ as giving a map from the cotangent space of the CB to the tangent plane of the fiber at $u$,
$\omega^{-1}:T^{*}_{u}\mathcal{C}\rightarrow T_{\mu}X_{u},$ (11)
where $\mu$ is a generic point on $X_{u}$, thus giving an isomorphism of vector spaces.
2.3 SK stratification of the CB
In extended supersymmetry, a non-zero SUSY central charge implies a non-trivial lower (BPS) bound on the masses, $M$, of charged states. In the case that we are analysing here, the central charge, for a given vacuum $u\in\mathcal{C}$ and charge $\mathbf{p}\in\Lambda_{u}$, is given by Witten:1978mh
$Z_{\mathbf{p}}(u)=\mathbf{p}^{T}\sigma(u).$ (12)
It follows from the BPS bound, $M\geq|Z|$, that states with charge $\mathbf{p}$ can only become massless at zeros of the locally holomorphic function $Z_{\mathbf{p}}(u)$. As this would result in massless charged states in the effective $\mathrm{U}(1)^{r}$ theory, the IR effective action, written in terms of free vector multiplets, breaks down, as reflected in non-analyticities of the CB metric. We assume all non-analyticities of the CB, $\mathcal{V}\subset\mathcal{C}$, have this form.555 Metric singularities at finite distance in moduli space that are not associated to the presence of extra particle-like massless degrees of freedom (corresponding to a vanishing central charge) would indicate the theory at hand has potentially more exotic massless degrees of freedom, such as a tensionless string. Non-analyticities of the metric along $\mathcal{V}$ imply $Z_{\mathbf{p}}$ is non-analytic there for those charges $\mathbf{p}$ corresponding to BPS states in the spectrum. This, in turn, implies non-analyticities of the special coordinates $\sigma$ of a special form at $\mathcal{V}$. Requiring the IR effective action to be physically consistent in the vicinity of $\mathcal{V}$ implies Argyres:2018urp :
  • $\mathcal{V}$ is closed in $\mathcal{C}$. If it were not, there would be no consistent physical interpretation of the IR effective action at the boundary points which are not contained within $\mathcal{V}$.
  • The Kähler metric extends over $\mathcal{V}$. It follows all distances on $\mathcal{V}$ are finite and well-defined.
  • As $Z_{\mathbf{p}}=0$ on $\mathcal{V}$, we can take $\mathcal{V}$ to be a union over components which vanish for a given value $\mathbf{p}$.
Further we assume $\mathcal{V}$ is a complex analytic set in $\mathcal{C}$. This is to ensure the non-existence of accumulation points in the complex plane transverse to $\mathcal{V}$. This assumption may not be necessary, as it might follow from the local holomorphicity of $\mathcal{V}$ and the fact that there are only a countably infinite number of central charges whose zeros can define $\mathcal{V}$ Argyres:2018urp . We can conclude from this that $\mathcal{V}$ is a complex co-dimension 1 variety in $\mathcal{C}$, and a generic point in $\mathcal{V}$ is a regular complex hypersurface in $\mathcal{C}$. We can write $\mathcal{V}$ as a union of its co-dimension 1 irreducible components,
$\mathcal{V}=\bigcup_{a}\ \mathcal{V}_{a}.$ (13)
The non-analyticity of the special section along one $\mathcal{V}_{a}$ component is reflected in there being a non-trivial EM monodromy, $M_{a}\in\mathrm{Sp}_{J}(2r,\mathbb{Z})$ around a path $\gamma_{a}\in\pi_{1}(\mathcal{C}\setminus\mathcal{V})$ linking only $\mathcal{V}_{a}$. The characterization of these possible linking monodromies, $M_{a}$, and their connection to rank-1 “transverse slice” SK geometries are described in more detail in Martone:2020nsy ; Argyres:2020wmq ; Cecotti:2021ouq .
But more interesting for our purpose is the fact Argyres:2020wmq ; Cecotti:2023ksl that the rank-$(r-1)$ $\mathcal{V}_{a}$ subvarieties inherit SK geometries of their own. These geometries, in turn, have co-dimension-1 singular components (i.e., where their inherited metrics have non-analyticities), and so forth, leading to an SK stratification of the CB. In terms of the $\mathcal{V}_{a}$ subvarieties, (the closure of) a co-dimension $s$ stratum of $\mathcal{C}$ is a connected component of an $s$-fold intersection of co-dimension-1 singular components, $\mathcal{V}_{a_{1}}\cap\cdots\cap\mathcal{V}_{a_{s}}$ (meant to include the cases where some of the intersections may be self-intersections).
This SK stratification can be understood in terms of the algebraically integrable system picture of SK geometry as follows Cecotti:2023ksl . According to the analysis of HO2007 ; HO2009 the singular fiber, $X_{u}$, at a regular point of a co-dimension-1 SK stratum, $u\in\mathcal{V}_{a}$, can be resolved into a set of transversely intersecting components, all of which are fiber bundles with a $\mathbb{P}^{1}$ fiber over a rank $r{-}1$ polarized abelian variety, $A_{u}$. The fibration of $A_{u}$ over $\mathcal{V}_{a}$ together with the restriction of the symplectic form is then a rank-$(r{-}1)$ SK geometry in its own right, thus giving the SK stratification.666 There are a number of caveats Cecotti:2023ksl . First, this works only with the assumption of the existence of a section of the abelian fibration over $\mathcal{C}$, which restricts the behavior of the singular fibers such that the symplectic reduction exists. The existence of a section is a requirement of a physical CB geometry, and is inherited by the SK stratification. Second, for certain “$I_{n}$-type” singular fibers the inherited SK geometry on the stratum is less constrained than those of physical CBs Argyres:2020wmq ; Cecotti:2021ouq . (These were called “irregular geometries” in Argyres:2020wmq .) Their existence does not affect the argument given here. Third, the analysis of HO2007 ; HO2009 assumes that the total space of the integrable system is a manifold, which is not obviously a requirement of a physical SK geometry. The analysis of Argyres:2020wmq , though less rigorous, indicates that the SK stratification nevertheless persists. Finally, it is not completely obvious that the singularities of the SK fibrations inherited by the strata are necessarily tame enough to continue the stratification down in dimension. Examples seem to support the hypothesis that the SK stratification continues “down”, but these examples are mostly at low rank. In particular, the restriction of the inverse symplectic form $\omega^{-1}$ to the resolved singular fiber (i.e., the $\mathbb{P}^{1}$ fiber bundle over $A_{u}$) still gives an isomorphism $T_{u}^{*}\mathcal{C}\cong T_{\mu}X_{u}$ as in (11), though now the tangent space to the (appropriate component of the resolved) singular fiber, $T_{\mu}X_{u}$, has direction tangent to its $A_{u}$ abelian variety base as well as to its $\mathbb{P}^{1}$ fiber. In particular, the $(r{-}1)$-dimensional span of the cotangents to the $\mathcal{V}_{a}$ stratum at $u$ map to the tangent space of the $A_{u}$ polarized abelian variety in the fiber.
Iterating this argument to lower dimensional strata, i.e., by restricting to higher-codimension subvarieties of the CB with additional metric non-analyticities, we find the following picture of the (resolved) singular fiber there. On an $\ell$-dimensional stratum, $\mathcal{V}_{\ell;a}$, the resolved singular fiber at a general point $u\in\mathcal{V}_{\ell;a}$ is a collection of intersecting components each of which are fiber bundles over a $-\ell$ abelian variety, $A^{\ell;a}_{u}$. And the tangent space to one of these fiber bundle components is isomorphic via $\omega^{-1}$ to the cotangent space of the base, with cotangents to the stratum mapped to tangents to $A^{\ell;a}_{u}$.
As the dilatations and $\mathrm{U}(1)_{R}$ rotations are symmetries, their $\mathbb{C}^{\times}$ action on the CB will, in particular, preserve the spectrum of charged states of the theory as well as the low energy effective action on the CB. This implies that it acts as an automorphism of the abelian variety fiber of the integrable system and preserves the symplectic form. That is, the holomorphic vector field $\mathcal{E}$ defined in (3) extends to one on the total space of the algebraically integrable system such that
$e^{z\mathcal{E}}\circ X_{u}$ $=X_{e^{z\mathcal{E}}u},$ $e^{z\mathcal{E}}\circ\omega$ $=e^{z}\omega.$ (14)
This last follows from the fact that weight of $\omega$ under the $\mathbb{C}^{\times}$ action on $\mathcal{C}$ is $\Delta_{\omega}=1$ since the fiber periods of $\omega$ are the special coordinates (10) which have mass dimension 1. The $\mathbb{C}^{\times}$ action therefore preserves each co-dimension-1 singular component, $\mathcal{V}_{a}$, as well as each of the SK strata defined by their intersections. That is to say, each $\ell$-dimensional SK stratum $\mathcal{V}_{\ell;a}$ is itself a union of orbits of the $\mathbb{C}^{\times}$ action.
3 Constraints on tuples from automorphisms of abelian varieties
In this section we will show how to use the $\mathbb{C}^{\times}$ symmetry action together with the SK stratification of the CB to put strong constraints on the spectrum of possible CB scaling dimensions. The key results are Properties 1 and 2, derived in the next subsection. These are closely related to earlier results described in Argyres:2018zay ; Caorsi:2018zsq ; Argyres:2018urp ; Argyres:2020wmq and Property 2 appears as “Fact 10” in section 2.12 of Cecotti:2023ksl . Then in subsection 3.2 we derive the algebraic constraints on compatible tuples of CB scaling dimensions which follow from these Properties.
3.1 Genuine rank-$r$ scaling dimensions
The set of scaling dimensions appearing at a given rank can be ordered by the smallest rank at which they first occur in a CB geometry. We call a scaling dimension a genuine $-\ell$ (or $\ell$-genuine) scaling dimension if it appears in CB geometries with rank $\ell$ and higher. [^7] These are referred to as new dimensions in $-\ell$ in Cecotti:2023ksl . This is a sensible notion since if a scaling dimension occurs at rank $\ell$, it necessarily occurs at all higher ranks, if only because higher-rank CB geometries can be formed by taking products of lower-rank geometries.
In general, the intersection of the $\mathbb{C}^{\times}$-strata $\mathcal{I}_{i_{1}\ldots i_{\ell}}$ defined in (5) with the SK strata $\mathcal{V}_{\ell^{\prime};a}$ may be very complicated. But the intersections of SK strata with the 1-dimensional $\mathbb{C}^{\times}$-strata $\mathcal{I}_{i}$ — the $u_{i}$ coordinate axes — are simple: since they are a single $\mathbb{C}^{\times}$ orbit, they either are wholly contained in a given SK stratum or do not intersect it at all. [^8] Technically, for this to be true, we are defining strata as the open sets (manifolds) formed by subtracting from their closure any proper sub-strata. A $\mathbb{C}^{\times}$ stratum $\mathcal{I}_{i_{1}\ldots i_{\ell}}$ comes with associated dimensions $(\Delta_{i_{1}},\ldots,\Delta_{i_{\ell}})$, which, recall, are the scaling dimensions of the CB scaling coordinates which are not set to zero on the stratum.
The key properties of the spectrum of CB branch scaling dimensions and the structure of CB singularities which follow from the $\mathbb{C}^{\times}$ symmetry and the SK stratification are:
Property 1.
$r$-genuine scaling dimensions belong to the set
[…]
Property 2.
If a rank-$r$ $\mathcal{N}{=}2$ SCFT has a CB operator with an $\ell$-genuine scaling dimension $\Delta_{i}$, the associated $u_{i}$ coordinate axis $\mathcal{I}_{i}$ is not contained in an SK stratum of dimension less than $\ell$.
We will now show how Properties 1 and 2 follow from the $\mathbb{C}^{\times}$ symmetry action on the CB and the SK stratification. Their derivations are closely related, and we will see that Property 2 is almost a corollary of Property 1.
First, recall that the $\mathbb{C}^{\times}$ action on the CB is given by (4) and on the abelian fibers of the integrable system by (14). If $\exp(z\mathcal{E})$ is in the stabilizer of $u\in\mathcal{C}$, i.e., if $\exp(z\mathcal{E})\cdot u=u$, then by (14) it defines an automorphism of the fiber, $\exp(z\mathcal{E})\cdot X_{u}\cong X_{u}$.
Now consider the $u_{i}$ coordinate axis, $\mathcal{I}_{i}$, and consider a point on it, $u\in\mathcal{I}_{i}$. Since $\mathcal{I}_{i}$ is a $\mathbb{C}^{\times}$ orbit, there is some value of $z$ such that $\exp(z\mathcal{E})$ fixes $u$. Indeed, from its action (4), it follows that
$\xi^{i}\doteq\exp\left(\frac{2\pi i}{\Delta_{i}}\mathcal{E}\right)$ (16)
generates the discrete subgroup of the $\mathrm{U}(1)_{R}$ symmetry which fixes $u$. (Indeed, it fixes all of $\mathcal{I}_{i}$ pointwise.) Furthermore it generates an automorphism of the fiber $X_{u}$. From the action (4) of $\mathcal{E}$ on the $\mathcal{C}$ (the CB), it follows that $\xi^{i}$ acts on the CB cotangent space by
$\mathop{}\!\mathrm{d}u_{j}$ $\mapsto\xi^{i}(\mathop{}\!\mathrm{d}u_{j})=\exp{\left(\frac{2\pi i}{\Delta_{i}}\Delta_{j}\right)}\mathop{}\!\mathrm{d}u_{j}.$ (17)
Also, even though the fiber $X_{u}$ might be singular (if $\mathcal{I}_{i}$ belongs to an SK stratum), an appropriate notion of $\omega^{-1}$ still exists, as explained in the last section. Furthermore, the (14) action implies that $\omega^{-1}$ transforms with weight $\Delta_{\omega^{-1}}=-1$, so $\xi^{i}$ acts on it as
$\omega^{-1}$ $\mapsto\xi^{i}(\omega^{-1})=\exp{\left(-\frac{2\pi i}{\Delta_{i}}\right)}\omega^{-1}.$ (18)
By composing the above actions, we obtain the corresponding action on the tangent space of the (appropriate component of the perhaps singular) fiber, $T_{\mu}X_{u}$,
$\omega^{-1}\circ\mathop{}\!\mathrm{d}u_{j}$ $\mapsto\xi^{i}(\omega^{-1})\circ\xi^{i}(\mathop{}\!\mathrm{d}u_{j})=\exp{\left(\frac{2\pi i(\Delta_{j}-1)}{\Delta_{i}}\right)}(\omega^{-1}\circ\mathop{}\!\mathrm{d}u_{j}),$ (19)
for $j=1,\ldots,r$. Denoting this representation by $\widehat{\rho}_{a}(\xi^{i})\in\mathrm{GL}(r,\mathbb{C})$, it has
$\text{eigenvalues of }\widehat{\rho}_{a}(\xi^{i})=\Bigl{\{}\exp\bigl{(}2\pi i(\Delta_{j}{-}1)/\Delta_{i}\bigr{)},\ \ j=1,\ldots,r\Bigr{\}}.$ (20)
Now suppose the $u_{i}$ coordinate axis $\mathcal{I}_{i}$ belongs to a stratum of singular fibers of some dimension $0<\ell\leq r$. (The case $\ell=r$ means $\mathcal{I}_{i}$ does not belong to any stratum of singularities, i.e., is regular.) By last section’s discussion of the SK stratification, the singular fibers of an $\ell$-dimensional stratum have an $\ell$ complex dimensional abelian variety factor, $A_{u}$, and $\omega^{-1}$ maps cotangent directions to the stratum to tangent directions of $A_{u}$. Since, by definition, $\mathop{}\!\mathrm{d}u_{i}$ is cotangent to the stratum containing $\mathcal{I}_{i}$, we see that $\xi^{i}$ acts as an automorphism of $A_{u}$ with eigenvalue $\exp(-2\pi i/\Delta_{i})$. Note that $\ell-1$ other values taken from the set (20) will also be eigenvalues of this automorphism, since the tangent space to $A_{u}$ is $\ell$-complex-dimensional. In other words, the above $\widehat{\rho}_{a}(\xi^{i})$ representation restricts to a representation
$\rho_{a}\doteq\widehat{\rho}_{a}\Bigl{|}_{A_{u}}:\mathrm{Aut}(A_{u})\to\mathrm{GL}(\ell,\mathbb{C}),$ (21)
of the automorphism group of the fiber abelian variety factor $A_{u}$, acting on its tangent space $T_{0}A_{u}$.
This fact places very strong constraints on the possible value of the scaling dimension $\Delta_{i}$ since the eigenvalues of automorphisms of abelian varieties can only belong to a small set of roots of unity. We will review this argument Argyres:2018urp ; Caorsi:2018zsq ; Argyres:2018zay now.
Note that any automorphism $\xi^{i}:A_{u}\to A_{u}$ of an abelian variety $A_{u}=\mathbb{C}^{\ell}/\Lambda$ can be uniquely lifted to a $\mathbb{C}$-linear map $\rho_{a}(\xi^{i})\in\mathrm{GL}(\ell,\mathbb{C})$ on the covering space (which is the tangent space) with the property $\rho_{a}(\xi^{i})(\Lambda)=\Lambda$. This gives the representation (21) of the automorphism group. On the other hand, we could equally consider the restriction $\rho_{r}(\xi^{i})=\rho_{a}(\xi^{i})\bigl{|}_{\Lambda}$ of $\rho_{a}$ to the lattice $\Lambda$ to get an integral representation
$\rho_{r}:\mathrm{Aut}(A_{u})\to\mathrm{Sp}_{J}(2\ell,\mathbb{Z}).$ (22)
The image has to be in $\mathrm{Sp}_{J}(2\ell,\mathbb{Z})$ since an automorphism preserves not only the lattice but also its polarization $J$. We call $\rho_{a}$ and $\rho_{r}$ the analytic representation and rational representation of $\mathrm{Aut}(A_{u})$, respectively. These two representations are related by
$\rho_{r}\otimes\mathbf{1}\cong\rho_{a}\oplus\bar{\rho_{a}},$ (23)
where $\mathbf{1}$ is the trivial one dimensional complex representation of $\mathrm{Aut}(A_{u})$ since if $A$, $R$ are matrices representing $\rho_{a}(\xi^{i})$, $\rho_{r}(\xi^{i})$, respectively, then they are related by the period matrix (8) by $A\Pi=\Pi R$. As $R$ is integer valued, complex conjugation gives $\bar{A}\,\bar{\Pi}=\bar{\Pi}R$, giving $\begin{pmatrix}A&0\\ 0&\bar{A}\end{pmatrix}\begin{pmatrix}\Pi\\ \bar{\Pi}\end{pmatrix}=\begin{pmatrix}\Pi\\ \bar{\Pi}\end{pmatrix}R$, which, by (9), is an isomorphism (23) between the two representations.
Returning to our setup, (23) tells us that, when accompanied by the conjugate representation, the analytic representation (21) is actually equivalent to an integral representation $\rho_{r}(\xi^{i})\in\mathrm{Sp}_{J}(2\ell,\mathbb{Z})$. So, in particular, $\exp(-2\pi i/\Delta_{i})$ must occur as the eigenvalue of an $\mathrm{Sp}_{J}(2\ell,\mathbb{Z})$ matrix all of whose eigenvalues are taken from the set (20). This is very constraining due to the fact that the characteristic polynomial of an $\mathrm{Sp}_{J}(2\ell,\mathbb{Z})$ matrix with unit norm eigenvalues can be written in the form
$\mathrm{char}(\rho_{r}(\xi^{i}))(z)$ $=\prod_{k\geq 1}\Phi_{k}(z)^{n_{k}},$ $\sum_{k\geq 1}\varphi(k)\,n_{k}$ $=2\ell,$ (24)
where $n_{k}$ are some non-negative integers, $\varphi$ is the Euler totient function, and $\Phi_{\ell}$ is the $k^{\text{th}}$ cyclotomic polynomial
$\Phi_{k}(z)$ $\doteq\!\!\!\prod_{\begin{subarray}{c}1\leq m\leq k\\ \gcd(k,m)=1\end{subarray}}\!\!\!(z-e^{2\pi im/k}).$ (25)
This follows because the characteristic polynomial of an integral matrix has integer coefficients, and the cyclotomic polynomials are the unique irreducible polynomials over the integers whose roots are roots of unity. The second condition in (24) comes from the fact that $\mathrm{deg}(\Phi_{k})=\varphi(k)$, and that $\mathrm{deg}(\mathrm{char}(\rho_{r}(\xi^{i})))=2\ell$. It implies that the only cyclotomic polynomials that can appear in $\mathrm{char}(\rho_{r}(\xi^{i}))$ are those with
$\varphi(k)\leq 2\ell,$ (26)
and so the eigenvalues of $\rho_{r}(\xi^{i})$ are in the set
$\text{eigenvalues of }\rho_{r}(\xi^{i})\in\Bigl{\{}\exp\bigl{(}2\pi im/k\bigr{)},\ 1\leq m\leq k,\ \gcd(k,m)=1,\ \varphi(k)\leq 2\ell\ \Bigr{\}}.$ (27)
Since $\exp(-2\pi i/\Delta_{i})$ is in this set, and since, by unitarity, $\Delta_{i}>1$, we learn that $\Delta_{i}$ must belong to the set (1) (with $r=\ell$). In particular, the possible new scaling dimensions at rank $\ell$ are those that saturate totient bound (26), giving Property 1. The result recorded in (15) is the subset of (1) with $\varphi(n)=2r$, where we have reparameterized it in terms of $d\doteq n$ and $a\doteq n-m$ for later convenience.
Now suppose $\Delta_{i}$ is an $\ell$-genuine scaling dimension, so it is an element of the set (15) with $r=\ell$. Then, if the $\mathcal{I}_{i}$ coordinate axis is contained in an SK stratum of dimension $k<\ell$, then the SK stratification implies $\exp(-2\pi i/\Delta_{i})$ is an eigenvalue of an automorphism of a dimension-$k$ abelian variety fiber, while the above argument implies that such a $\Delta_{i}$ in the set (1) with $r=k$, which does not include (15) with $r=\ell$. This contradiction shows that $\mathcal{I}_{i}$ cannot be contained in any SK stratum of dimension $k<\ell$, thus showing Property 2.
Immediate consequences of Property 2 are
Property 3.
If a rank-$r$ $\mathcal{N}{=}2$ SCFT has a CB operator with an $r$-genuine scaling dimension $\Delta_{i}$, the associated $u_{i}$ coordinate axis $\mathcal{I}_{i}$ is necessarily non-singular,
and, with a bit more work, its generalization,
Property 4.
If an $\ell$-dimensional $\mathbb{C}^{\times}$-stratum, $\mathcal{I}_{i_{1}\ldots i_{\ell}}$, is contained in an $\ell$-dimensional SK stratum, then its associated scaling dimensions are all allowed at rank $\ell$.
This latter follows from the fact that if $\mathcal{I}_{i_{1}\ldots i_{\ell}}$ is contained in an $\ell$-dimensional SK stratum, then a subset of $\ell$ of the eigenvalues in (20) will be eigenvalues of an $\mathrm{Sp}_{J}(2\ell,\mathbb{Z})$ matrix.
These properties imply that subsets of $r$-tuples of scaling dimensions must satisfy the consistency conditions. Define a genuine $\ell$-tuple to be an $\ell$-tuple of $k$-genuine scaling dimensions with $k\leq\ell$ and with a least one $\ell$-genuine scaling dimension saturating this inequality. Then Properties 14 imply
Property 5.
If $\mathcal{D}=\{\Delta_{1},\ldots,\Delta_{r}\}$ is an $r$-tuple of scaling dimensions of a rank-$r$ CB, and $\Delta_{i}$ is genuine in rank-$\ell_{i}$, then there is a genuine $\ell_{i}$-tuple contained in $\mathcal{D}$ that also contains $\Delta_{i}$.
Aside.
Note that for certain ranks $r$, there are no $r$-genuine scaling dimensions, i.e., the set (15) may be empty. This is due to the existence of non-totient numbers; that is, numbers $q$ for which no solution to $\varphi(p)=q$ exists. Of course, all odd numbers greater than $1$ are non-totient, but there are, in fact, infinitely many even non-totient numbers too. The first few are (A005277 in OEIS):
14, 26, 34, 38, 50, 62, 68, 74, 76, 86, 90,.1426343850626874768690 14,\ 26,\ 34,\ 38,\ 50,\ 62,\ 68,\ 74,\ 76,\ 86,\ 90$\ldots$.14 , 26 , 34 , 38 , 50 , 62 , 68 , 74 , 76 , 86 , 90 , … . (28)
Comparing with (15), we see that SCFTs of rank-$7$ have no genuinely rank-$7$ scaling dimensions, for example. In these cases, a non-singular $\mathbb{C}^{\times}$ stratum in the CB need not exist.
3.2 Constraints on $r$-genuine tuples
We now use the properties derived in the previous subsection to constrain the sets of allowed $r$-tuples of CB scaling dimensions. In this subsection we will focus on genuine $r$-tuples at rank $r$. Recall that these are tuples of scaling dimensions of a rank-$r$ CB with at least one genuine rank-$r$ scaling dimension. Then, in the next subsection we will show how the non-genuine tuples at a given rank can be formed from genuine tuples of lower rank.
Since we have assumed that the given $r$-tuple is genuine, there is always (at least) one scaling dimension that is genuine; call it $\Delta_{i}$, and take $u\in\mathcal{I}_{i}$. By property 3, the $u_{i}$-coordinate axis, $\mathcal{I}_{i}$, is necessarily non-singular. Since $\mathcal{I}_{i}$ is non-singular, its abelian fiber has dimension $r$, so $\widehat{\rho}_{a}(\xi^{i})$ is the analytic representation of an automorphism of a dimension-$r$ abelian variety. Thus, all its eigenvalues, (20), and by (23) their conjugates as well, are the eigenvalues of an $\mathrm{Sp}_{J}(2r,\mathbb{Z})$ matrix. Thus the roots of its characteristic polynomial are exactly
$\lambda_{j}^{\pm}$ $=\exp\left(\pm 2\pi i\frac{\Delta_{j}-1}{\Delta_{i}}\right),$ $j$ $=1,\ldots,r.$ (29)
Now recall that $\Delta_{i}$ is a genuinely rank-$r$ scaling dimension. Thus we can write, as in (15), $\Delta_{i}=d_{i}/(d_{i}-a_{i})$ for some coprime integers $a_{i}$ and $d_{i}$ satisfying $a_{i}\leq d_{i}-1$ and $\varphi(d_{i})=2r$. As such,
$\lambda_{i}^{+}=\exp\left(-2\pi i\frac{d_{i}-a_{i}}{d_{i}}\right),$ (30)
is a root of $\mathrm{char}(\rho_{r}(\xi^{i}))$. Since this is a root of the cyclotomic polynomial $\Phi_{d_{i}}$, and since this polynomial has $2r$ distinct roots (the primitive $d_{i}$th roots of unity), we must have that the characteristic polynomial (24) is in fact $\mathrm{char}(\rho_{r}(\xi^{i}))=\Phi_{d_{i}}$. By comparing with the general form of a cyclotomic polynomial, we get that the other roots give rise to the relations
$\frac{\Delta_{j}-1}{\Delta_{i}}=\frac{a_{j}}{d_{i}}\ \ \Longrightarrow\ \ \Delta_{j}=1+\frac{a_{j}}{d_{i}-a_{i}},\qquad\forall j\neq i,$ (31)
for some other integers $a_{j}$ coprime to $d_{i}$.
Notice that (31) must be satisfied for all genuine scaling dimensions $\Delta_{i}$ against all other scaling dimensions $\Delta_{j}$ in the candidate $r$-tuple. The set of relations (31) may be satisfied for one genuine scaling dimension $\Delta_{i}$, but the corresponding set of relations may fail to be satisfied relative to a different genuine scaling dimensions in the $r$-tuple (should one exist). We call this feature the non-reflexivity of the genuine scaling dimension condition — see, e.g., example 3.2 below.
The fact that the roots of $\mathrm{char}(\rho_{r}(\xi^{i}))=\Phi_{d_{i}}$ are all distinct has several non-trivial consequences, which we now outline.
  1. 1.
    Since each $\lambda^{\pm}_{j}$ must correspond to a different root of unity, we get the modular constraint
    $a_{j}-a_{k}\neq 0\quad(\mathrm{mod}\,d_{i}),\quad\forall j\neq k.$ (32)
  2. 2.
    The eigenvalues of an $\mathrm{Sp}_{J}(2r,\mathbb{Z})$ matrix come in reciprocal pairs $\{\lambda_{j}^{+},\lambda_{j}^{-}\}$. These reciprocal pairs signal the existence of a single CB scaling dimension, so we must take care to not count these as separate putative CB dimensions. This gives the constraint
    $a_{j}+a_{k}\neq 0\quad(\mathrm{mod}\,d_{i}),\quad\forall j\neq k.$ (33)
  3. 3.
    An immediate consequence of the previous points is that the presence of an $r$-genuine CB dimension forbids repeated dimensions. Repeated dimensions can, however, occur in non-genuinely rank-$r$ tuples.
  4. 4.
    Another interesting corollary of (33) is that we cannot have a CB scaling dimension $\Delta_{j}=2$ in genuine $r$-tuples. Indeed, this would correspond to the eigenvalue
    $\lambda_{i}^{-}=\exp\left(2\pi i\frac{1}{\,\Delta_{i}}\right),$ (34)
    but this is the reciprocal of the eigenvalue corresponding to $\Delta_{i}$, which is genuinely rank-$r$, leading to a contradiction.
  5. 5.
    Also, if we take $\Delta_{i}$ to be an integer genuinely rank-$r$ scaling dimension, equation (31) tells us that all other dimensions are also integer. Furthermore, if $\Delta_{i}$ is even, then so are the other compatible dimensions.
To summarize, equations (31), (32), and (33) are our main results constraining the possible sets of genuinely rank-$r$ tuples.
To illustrate these constraints, let us now consider some low rank examples.
Example 3.1.
Consider the genuinely rank-$2$ scaling dimension $\Delta_{1}=\tfrac{8}{7}$. In order for another CB scaling dimension $\Delta_{2}$ to be consistent, we must have
$\Delta_{2}=1+\frac{a_{2}}{7},\ \ \gcd(a_{2},8)=1.$ (35)
[…]
Example 3.2.
Consider the genuinely rank-$3$ scaling dimension $\Delta_{1}=\tfrac{18}{17}$, so $d_{1}=18$ and $a_{1}=1$. The consistent triplets of CB scaling dimensions $\{\Delta_{1},\Delta_{2},\Delta_{3}\}$ must satisfy
$\Delta_{j}=1+\frac{a_{j}}{17},\ \ \gcd(a_{j},18)=1,\ \ j\in\{2,3\}.$ (36)
[…]
3.3 Constraints on non-genuine $r$-tuples.
Key to our discussion in the genuine case was the fact that the $u_{i}$-plane $\mathcal{I}_{i}$ was non-singular. This can no longer be guaranteed in the non-genuine case, as any $(r-1)$-subtuple is allowed at rank-$(r-1)$. Nevertheless, the scaling dimensions are still constrained by Property 5, derived above.
This property implies that a non-genuine $r$-tuple is given as a union of genuine $\ell$-tuples for $\ell<r$. But we have to be careful about what is meant by “union” here. In particular, this means that we can obtain a non-genuine $r$-tuple by combining genuine tuples of lower rank up to overlapping dimensions, or by repeating entries in a lower-rank genuine tuple. For example, if $\{\Delta_{1},\Delta_{2}\}$ is a genuinely rank-2 tuple, then $\{\Delta_{1},\Delta_{2},\Delta_{2}\}$ is an allowed non-genuine rank-3 tuple.
In the next section we outline a systematic procedure to generate all allowed $r$-tuples, whether genuine or not.
4 Allowed $r$-tuples of CB scaling dimensions
We present an algorithm for constructing the allowed $r$-tuples of CB operator scaling dimensions at a given rank $r$. Our algorithm considers the genuine rank-$r$ tuples, and non-genuine rank-$r$ tuples separately.
Genuine $r$-tuples.
To calculate the allowed genuinely rank-$r$ tuples, we generalize the procedures outlined in examples 3.1 and 3.2.
  1. 1.
    Select a genuine rank-$r$ scaling dimension.
  2. 2.
    Using (31) and the relations between the $\{a_{j}\}$ calculate which of the elements of (1) can occupy a tuple with the chosen genuine dimension.
  3. 3.
    Generate all possible $r$-tuples of the genuine dimension with the dimensions it can occupy a tuple with.
  4. 4.
    Note that if a tuple contains more than 1 genuinely rank-$r$ dimension, then we must check our consistency conditions with all of them to ensure the tuple is consistent.
Repeating this process for all of the genuine rank-$r$ scaling dimensions completely determines the set of genuinely rank-$r$ tuples. In appendix A we present the results of this algorithm for rank-$2$, $3$ and $4$.
Non-genuine $r$-tuples.
These $r$-tuples are constructed from the lower rank tuples. Therefore, to calculate the non-genuine rank-$r$ tuples one generates all of the possible unordered combinations of lower-rank tuples whose ranks sum to $r$. One can allow for duplicates, as long as the total number of scaling dimensions is $r$. For example,
  • For rank-$2$ theories, the only option of a non-genuine pair is to take two rank-$1$ scaling dimensions, to form a rank-$2$ pair.
  • For rank-$3$ theories, we can not only take three rank-$1$ scaling dimensions or a rank-$2$ pair and a rank-$1$ dimension, but also two rank-$2$ pairs as long as they have at least one dimension in common so that there are 3 dimensions in total. An example is the rank-3 $(A_{1},D_{7})$ AD-theory which has scaling dimensions $\{\tfrac{8}{7},\tfrac{10}{7},\tfrac{12}{7}\}$. This is a non-genuine triple that arises as the combination of two genuine pairs $\{\tfrac{8}{7},\tfrac{10}{7}\}$ and $\{\tfrac{8}{7},\tfrac{12}{7}\}$, with one overlapping dimension $\tfrac{8}{7}$.
  • For rank-$4$ theories, proceed similarly. We can have a rank-$1$ scaling dimension combined with a choice of an allowed genuine or non-genuine rank-$3$ tuple, or all possible pairs of allowed genuine or non-genuine rank-$2$ tuples. Moreover with overlaps we can have two rank-$3$ triples (as long as they have two dimensions in common), or one rank-$3$ triple and one rank-$2$ pair (as long as they have one dimension in common), or three rank-$2$ pairs (such that there are 4 dimensions in total, e.g. three pairs of the form $\{\Delta_{1},\Delta_{2}\},\{\Delta_{2},\Delta_{3}\},\{\Delta_{3},\Delta_{4}\}$).
More generally, allowed non-genuine $r$-tuples can be obtained from combinations of allowed genuine and non-genuine $m$-tuples with $m<r$ algorithmically, by combining all possible ways of forming $r$ distinct scaling dimensions allowing all possible overlaps. The list at rank $r$ is obtained as follows
  • A choice of one allowed genuine $(r-1)$-tuple, and either a choice of an allowed rank-$1$ scaling dimension or a choice of an allowed collection of $m_{i}$-tuples with $1<m_{i}<r$ such that together with the scaling dimensions in the chosen $(r-1)$-tuple there are $r$ inequivalent scaling dimensions in total up to overlaps;
  • A choice of one allowed genuine $(r-2)$-tuple, and either a choice an allowed (genuine or non-genuine) rank-$2$ scaling dimension or a choice of an allowed collection of $m_{i}$-tuples with $2<m_{i}<r$ such that together with the scaling dimensions in the chosen $(r-2)$-tuple there are $r$ inequivalent scaling dimensions in total up to overlaps;
  • $\cdots$
  • The choice a genuine $(r-k)$-tuple and then either a genuine or non-genuine $k$-tuple, or all possible collections of genuine or non-genuine $m_{i}$-tuples with $k<m_{i}<r$ that together with the scaling dimensions in the chosen genuine $(r-k)$-tuple give rise to a collection of $r$ distinct scaling dimensions;
  • $\cdots$
  • A choice of one allowed genuine rank-$2$ scaling dimension, and either a choice of an allowed (genuine or non-genuine) rank-$(r-2)$ scaling dimension or a choice of an allowed collection of $m_{i}$-tuples with $(r-2)<m_{i}<r$ such that there are $r$ inequivalent scaling dimensions in total up to overlaps;
  • A choice of one allowed rank-$1$ scaling dimension, and either a choice of an allowed non-genuine rank-$(r-1)$ scaling dimension
Of course, the above algorithm generates all allowed non-genuine $r$-tuples, but it is still possible there is a finer classification: we expect that the collection of actually realized scaling dimensions in 4d $\mathcal{N}=2$ SCFTs is a subset of the list generated by the algorithm above.
Comparison with previous estimates.
To round out this section, let us compare our new estimates for the number of rank-$r$ tuples with the naïve estimates using the results of Caorsi:2018zsq ; Argyres:2018urp . As the rank grows large, the number of allowable rank-$r$ dimensions scales asymptotically as Caorsi:2018zsq
$N_{r}=\frac{2\zeta(2)\zeta(3)}{\zeta(6)}r^{2}+o(r^{2}).$ (37)
If we assume that any combination of scaling dimensions constitutes a valid rank-$r$ tuple, a standard stars and bars argument gives the naïve number of distinct $r$-tuples
$T_{r}^{\star}={N_{r}+r-1\choose r}\sim\mathcal{O}\left(\frac{r^{2r}}{r!}\right).$ (38)
Included in this count is the approximate number of non-genuine tuples given by
$L^{\star}_{r}={N_{r-1}+r-1\choose r}.$ (39)
As such, the difference $T_{r}^{\star}-L_{r}^{\star}$ gives a rudimentary counting of the number of genuine $r$-tuples. By imposing our conditions on this putative set of scaling dimensions, we can see that this is a vast overestimate. For ranks less than or equal to $4$, we tabulate our results in table 1.
rank $N_{r}$ $T_{r}^{\star}$ $L_{r}^{\star}$ $T_{r}^{\mathrm{gen}}$ $Q_{r}$
$2$ $23$ $276$ $28$ $35$ $1.41129\times 10^{-1}$
$3$ $47$ $18424$ $2300$ $242$ $1.50087\times 10^{-2}$
$4$ $87$ $2555190$ $230300$ $911$ $3.91846\times 10^{-4}$
Table 1: A comparison of our estimates of the number of genuinely rank-$r$ tuples to the naïve counting. Here $T_{r}^{\mathrm{gen}}$ is the number of genuinely rank-$r$ tuples consistent with our constraints and $Q_{r}=T_{r}^{\mathrm{gen}}/(T_{r}^{\star}-L_{r}^{\star})$ is the ratio of our estimates by the previous estimate on the number of genuine $r$-tuples.
It would be useful and informative to obtain precise bounds on the growth of the number of allowed $r$-tuples, and on the statistics of their distribution. See, for instance, Perlmutter:2024noo for an example of an interesting application of this kind of information.
Acknowledgements.
PCA is supported in part by DOE grant DE-SC1019775. The work of MDZ has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 851931), the VR project grant No. 2023-05590 and from the Simons Foundation (grant #888984, Simons Collaboration on Global Categorical Symmetries). MDZ also acknowledges the VR Centre for Geometry and Physics (VR grant No. 2022-06593). The work of BS and MM is supported by STFC grant ST/T000759/1. RM is supported by a Knut and Alice Wallenberg Foundation postdoctoral scholarship in mathematics.
Appendix A Tables of genuine $r$-tuples
In table 2 we present the valid genuinely rank-2 pairs. This corrects the findings of Kaidi:2022sng where 16 additional pairs were reported. These additional 16 pairs had two genuinely rank-2 scaling dimensions that are ruled out by checking our compatibility conditions using both of the scaling dimensions. For example, the pair $\left\{\tfrac{12}{5},8\right\}$ looks valid if we only check the compatibilty conditions for $\tfrac{12}{5}$ but, as mentioned in section 3.2, the only dimensions valid with $8$ are even integers. As such, this is an invalid pair.
We present the genuinely rank-3 triplets in tables 3 and 4, and the genuinely rank-4 quadruplets in tables 512. For tuples that contain more than one genuinely rank-$r$ scaling dimension, we have highlighted them in blue if they have already appeared in the table for a smaller scaling dimension $\Delta_{\mathrm{new}}$.
$d$$\{\Delta_{1},\Delta_{2}\}$
5 $\left\{\tfrac{5}{4},\tfrac{3}{2}\right\}$, $\left\{\tfrac{5}{4},3\right\}$, $\left\{\tfrac{5}{4},4\right\}$, $\left\{\frac{4}{3},\tfrac{5}{3}\right\}$, $\left\{\tfrac{5}{3},3\right\}$, $\left\{\tfrac{5}{3},4\right\}$, $\left\{\tfrac{3}{2},\tfrac{5}{2}\right\}$,
$\left\{\tfrac{5}{2},3\right\}$, $\left\{\tfrac{5}{2},4\right\}$, $\left\{3,5\right\}$, $\{4,5\}$
8 $\left\{\tfrac{8}{7},\tfrac{10}{7}\right\}$, $\left\{\tfrac{8}{7},\tfrac{12}{7}\right\}$, $\left\{\tfrac{8}{7},4\right\}$, $\left\{\tfrac{8}{7},6\right\}$, $\left\{\tfrac{6}{5},\tfrac{8}{5}\right\}$, $\left\{\tfrac{8}{5},4\right\}$, $\left\{\tfrac{8}{5},6\right\}$
$\left\{\tfrac{4}{3},\tfrac{8}{3}\right\}$, $\left\{\tfrac{8}{3},4\right\}$, $\left\{\tfrac{8}{3},6\right\}$, $\{4,8\}$, $\{6,8\},\{8,12\}$
10 $\left\{\tfrac{10}{9},\tfrac{4}{3}\right\}$, $\left\{\tfrac{10}{9},4\right\}$, $\left\{\tfrac{10}{7},4\right\}$, $\left\{\tfrac{4}{3},\tfrac{10}{3}\right\}$, $\left\{\tfrac{10}{3},4\right\}$, $\{4,10\}$
12 $\left\{\tfrac{12}{11},6\right\}$, $\left\{\tfrac{12}{7},6\right\}$, $\left\{\tfrac{6}{5},\tfrac{12}{5}\right\}$, $\left\{\tfrac{12}{5},6\right\}$, $\{6,12\}$
Table 2: The 35 genuine rank-$2$ scaling dimension pairs.
$\Delta_{\mathrm{new}}$$\{\Delta_{1},\Delta_{2},\Delta_{3}\}$
$\tfrac{18}{17}$$\left\{\tfrac{18}{17},6,12\right\},\left\{\tfrac{18}{17},6,8\right\}$
$\tfrac{14}{13}$$\left\{\tfrac{14}{13},10,12\right\},\left\{\tfrac{14}{13},6,12\right\},\left\{\tfrac{14}{13},4,10\right\},\left\{\tfrac{14}{13},4,6\right\}$
$\tfrac{9}{8}$$\left\{\tfrac{9}{8},\tfrac{5}{4},\tfrac{3}{2}\right\},\left\{\tfrac{9}{8},\tfrac{5}{4},6\right\},\left\{\tfrac{9}{8},\tfrac{5}{4},5\right\},\left\{\tfrac{9}{8},\tfrac{3}{2},8\right\},\left\{\tfrac{9}{8},6,8\right\},\left\{\tfrac{9}{8},5,8\right\}$
$\left\{\tfrac{9}{8},\tfrac{3}{2},3\right\},\left\{\tfrac{9}{8},\tfrac{3}{2},12\right\},\left\{\tfrac{9}{8},3,6\right\},\left\{\tfrac{9}{8},6,12\right\},\left\{\tfrac{9}{8},3,5\right\},\left\{\tfrac{9}{8},5,12\right\}$
$\tfrac{7}{6}$$\left\{\tfrac{7}{6},\tfrac{4}{3},\tfrac{3}{2}\right\},\left\{\tfrac{7}{6},\tfrac{4}{3},\tfrac{8}{3}\right\},\left\{\tfrac{7}{6},\tfrac{4}{3},5\right\},\left\{\tfrac{7}{6},\tfrac{4}{3},12\right\},\left\{\tfrac{7}{6},\tfrac{4}{3},\tfrac{5}{3}\right\},\left\{\tfrac{7}{6},\tfrac{4}{3},4\right\},\left\{\tfrac{7}{6},\tfrac{3}{2},\tfrac{5}{2}\right\},\left\{\tfrac{7}{6},\tfrac{5}{2},\tfrac{8}{3}\right\},\left\{\tfrac{7}{6},\tfrac{5}{2},5\right\},\left\{\tfrac{7}{6},\tfrac{5}{2},12\right\}$
$\left\{\tfrac{7}{6},\tfrac{5}{3},\tfrac{5}{2}\right\},\left\{\tfrac{7}{6},\tfrac{5}{2},4\right\},\left\{\tfrac{7}{6},\tfrac{3}{2},6\right\},\left\{\tfrac{7}{6},\tfrac{8}{3},6\right\},\left\{\tfrac{7}{6},5,6\right\},\left\{\tfrac{7}{6},6,12\right\},\left\{\tfrac{7}{6},\tfrac{5}{3},6\right\},\left\{\tfrac{7}{6},4,6\right\},\left\{\tfrac{7}{6},\tfrac{3}{2},3\right\},\left\{\tfrac{7}{6},\tfrac{3}{2},10\right\}$
$\left\{\tfrac{7}{6},\tfrac{8}{3},3\right\},\left\{\tfrac{7}{6},\tfrac{8}{3},10\right\},\left\{\tfrac{7}{6},3,5\right\},\left\{\tfrac{7}{6},5,10\right\},\left\{\tfrac{7}{6},3,12\right\},\left\{\tfrac{7}{6},10,12\right\},\left\{\tfrac{7}{6},\tfrac{5}{3},3\right\},\left\{\tfrac{7}{6},\tfrac{5}{3},10\right\},\left\{\tfrac{7}{6},3,4\right\},\left\{\tfrac{7}{6},4,10\right\}$
$\tfrac{14}{11}$$\left\{\tfrac{12}{11},\tfrac{14}{11},4\right\},\left\{\tfrac{12}{11},\tfrac{14}{11},12\right\},\left\{\tfrac{14}{11},4,10\right\},\left\{\tfrac{14}{11},10,12\right\},\left\{\tfrac{14}{11},4,6\right\},\left\{\tfrac{14}{11},6,12\right\}$
$\tfrac{9}{7}$$\left\{\tfrac{8}{7},\tfrac{9}{7},8\right\},\left\{\tfrac{8}{7},\tfrac{9}{7},\tfrac{12}{7}\right\},\left\{\tfrac{8}{7},\tfrac{9}{7},3\right\},\left\{\tfrac{8}{7},\tfrac{9}{7},12\right\},\left\{\tfrac{9}{7},5,8\right\},\left\{\tfrac{9}{7},\tfrac{12}{7},5\right\}$
$\left\{\tfrac{9}{7},3,5\right\},\left\{\tfrac{9}{7},5,12\right\},\left\{\tfrac{9}{7},6,8\right\},\left\{\tfrac{9}{7},\tfrac{12}{7},6\right\},\left\{\tfrac{9}{7},3,6\right\},\left\{\tfrac{9}{7},6,12\right\}$
$\tfrac{18}{13}$$\left\{\tfrac{18}{13},6,8\right\},\left\{\tfrac{18}{13},6,12\right\}$
$\tfrac{7}{5}$$\left\{\tfrac{6}{5},\tfrac{7}{5},\tfrac{8}{5}\right\},\left\{\tfrac{6}{5},\tfrac{7}{5},3\right\},\left\{\tfrac{6}{5},\tfrac{7}{5},10\right\},\left\{\tfrac{6}{5},\tfrac{7}{5},\tfrac{9}{5}\right\},\left\{\tfrac{6}{5},\tfrac{7}{5},6\right\},\left\{\tfrac{7}{5},\tfrac{8}{5},4\right\},\left\{\tfrac{7}{5},3,4\right\},\left\{\tfrac{7}{5},4,10\right\},\left\{\tfrac{7}{5},4,6\right\}$
$\left\{\tfrac{7}{5},\tfrac{8}{5},5\right\},\left\{\tfrac{7}{5},\tfrac{8}{5},12\right\},\left\{\tfrac{7}{5},3,5\right\},\left\{\tfrac{7}{5},3,12\right\},\left\{\tfrac{7}{5},5,10\right\},\left\{\tfrac{7}{5},10,12\right\},\left\{\tfrac{7}{5},\tfrac{9}{5},12\right\},\left\{\tfrac{7}{5},5,6\right\},\left\{\tfrac{7}{5},6,12\right\}$
$\tfrac{14}{9}$$\left\{\tfrac{10}{9},\tfrac{4}{3},\tfrac{14}{9}\right\},\left\{\tfrac{10}{9},\tfrac{14}{9},6\right\},\left\{\tfrac{10}{9},\tfrac{14}{9},10\right\},\left\{\tfrac{4}{3},\tfrac{14}{9},\tfrac{8}{3}\right\},\left\{\tfrac{14}{9},\tfrac{8}{3},6\right\},\left\{\tfrac{14}{9},\tfrac{8}{3},10\right\}$
$\left\{\tfrac{4}{3},\tfrac{14}{9},12\right\},\left\{\tfrac{14}{9},6,12\right\},\left\{\tfrac{14}{9},10,12\right\},\left\{\tfrac{4}{3},\tfrac{14}{9},4\right\},\left\{\tfrac{14}{9},4,6\right\},\left\{\tfrac{14}{9},4,10\right\}$
$\tfrac{18}{11}$$\left\{\tfrac{12}{11},\tfrac{18}{11},8\right\},\left\{\tfrac{12}{11},\tfrac{18}{11},12\right\},\left\{\tfrac{18}{11},6,8\right\},\left\{\tfrac{18}{11},6,12\right\}$
Table 3: The genuinely rank-$3$ sets of scaling dimensions with $\tfrac{18}{17}\leq\Delta_{\mathrm{new}}\leq\tfrac{18}{11}$.
$\Delta_{\mathrm{new}}$$\{\Delta_{1},\Delta_{2},\Delta_{3}\}$
$\tfrac{7}{4}$$\left\{\tfrac{5}{4},\tfrac{3}{2},\tfrac{7}{4}\right\},\left\{\tfrac{5}{4},\tfrac{7}{4},5\right\},\left\{\tfrac{5}{4},\tfrac{7}{4},12\right\},\left\{\tfrac{5}{4},\tfrac{7}{4},4\right\},\left\{\tfrac{3}{2},\tfrac{7}{4},3\right\},\left\{\tfrac{7}{4},3,5\right\},\left\{\tfrac{7}{4},3,12\right\},\left\{\tfrac{7}{4},3,4\right\},\left\{\tfrac{3}{2},\tfrac{7}{4},10\right\},\left\{\tfrac{7}{4},5,10\right\}$
$\left\{\tfrac{7}{4},10,12\right\},\left\{\tfrac{7}{4},4,10\right\},\left\{\tfrac{3}{2},\tfrac{7}{4},\tfrac{5}{2}\right\},\left\{\tfrac{3}{2},\tfrac{7}{4},6\right\},\left\{\tfrac{7}{4},\tfrac{5}{2},5\right\},\left\{\tfrac{7}{4},5,6\right\},\left\{\tfrac{7}{4},\tfrac{5}{2},12\right\},\left\{\tfrac{7}{4},6,12\right\},\left\{\tfrac{7}{4},\tfrac{5}{2},4\right\},\left\{\tfrac{7}{4},4,6\right\}$
$\tfrac{9}{5}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{6}{5},\tfrac{7}{5},\tfrac{9}{5}\right\}}},\left\{\tfrac{6}{5},\tfrac{9}{5},5\right\},\left\{\tfrac{6}{5},\tfrac{9}{5},\tfrac{12}{5}\right\},\left\{\tfrac{6}{5},\tfrac{9}{5},6\right\},\left\{\tfrac{9}{5},3,5\right\},\left\{\tfrac{9}{5},\tfrac{12}{5},3\right\},\left\{\tfrac{9}{5},3,6\right\}$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{7}{5},\tfrac{9}{5},12\right\}}},\left\{\tfrac{9}{5},5,12\right\},\left\{\tfrac{9}{5},\tfrac{12}{5},12\right\},\left\{\tfrac{9}{5},6,12\right\},\left\{\tfrac{9}{5},5,8\right\},\left\{\tfrac{9}{5},\tfrac{12}{5},8\right\},\left\{\tfrac{9}{5},6,8\right\}$
$\tfrac{9}{4}$$\left\{\tfrac{5}{4},\tfrac{3}{2},\tfrac{9}{4}\right\},\left\{\tfrac{5}{4},\tfrac{9}{4},6\right\},\left\{\tfrac{5}{4},\tfrac{9}{4},5\right\},\left\{\tfrac{3}{2},\tfrac{9}{4},8\right\},\left\{\tfrac{9}{4},6,8\right\},\left\{\tfrac{9}{4},5,8\right\}$
$\left\{\tfrac{3}{2},\tfrac{9}{4},3\right\},\left\{\tfrac{3}{2},\tfrac{9}{4},12\right\},\left\{\tfrac{9}{4},3,6\right\},\left\{\tfrac{9}{4},6,12\right\},\left\{\tfrac{9}{4},3,5\right\},\left\{\tfrac{9}{4},5,12\right\}$
$\tfrac{7}{3}$$\left\{\tfrac{4}{3},\tfrac{5}{3},\tfrac{7}{3}\right\},\left\{\tfrac{4}{3},\tfrac{7}{3},4\right\},\left\{\tfrac{4}{3},\tfrac{7}{3},\tfrac{8}{3}\right\},\left\{\tfrac{4}{3},\tfrac{7}{3},5\right\},\left\{\tfrac{4}{3},\tfrac{7}{3},12\right\},\left\{\tfrac{5}{3},\tfrac{7}{3},6\right\},\left\{\tfrac{7}{3},4,6\right\},\left\{\tfrac{7}{3},\tfrac{8}{3},6\right\},\left\{\tfrac{7}{3},5,6\right\},\left\{\tfrac{7}{3},6,12\right\}$
$\left\{\tfrac{5}{3},\tfrac{7}{3},3\right\},\left\{\tfrac{5}{3},\tfrac{7}{3},10\right\},\left\{\tfrac{7}{3},3,4\right\},\left\{\tfrac{7}{3},4,10\right\},\left\{\tfrac{7}{3},\tfrac{8}{3},3\right\},\left\{\tfrac{7}{3},\tfrac{8}{3},10\right\},\left\{\tfrac{7}{3},3,5\right\},\left\{\tfrac{7}{3},5,10\right\},\left\{\tfrac{7}{3},3,12\right\},\left\{\tfrac{7}{3},10,12\right\}$
$\tfrac{18}{7}$$\left\{\tfrac{8}{7},\tfrac{12}{7},\tfrac{18}{7}\right\},\left\{\tfrac{8}{7},\tfrac{18}{7},12\right\},\left\{\tfrac{8}{7},\tfrac{18}{7},8\right\},\left\{\tfrac{12}{7},\tfrac{18}{7},6\right\},\left\{\tfrac{18}{7},6,12\right\},\left\{\tfrac{18}{7},6,8\right\}$
$\tfrac{14}{5}$$\left\{\tfrac{6}{5},\tfrac{8}{5},\tfrac{14}{5}\right\},\left\{\tfrac{6}{5},\tfrac{14}{5},10\right\},\left\{\tfrac{6}{5},\tfrac{14}{5},6\right\},\left\{\tfrac{8}{5},\tfrac{14}{5},4\right\},\left\{\tfrac{14}{5},4,10\right\},\left\{\tfrac{14}{5},4,6\right\},\left\{\tfrac{8}{5},\tfrac{14}{5},12\right\},\left\{\tfrac{14}{5},10,12\right\},\left\{\tfrac{14}{5},6,12\right\}$
$\tfrac{7}{2}$$\left\{\tfrac{3}{2},\tfrac{5}{2},\tfrac{7}{2}\right\},\left\{\tfrac{3}{2},\tfrac{7}{2},6\right\},\left\{\tfrac{3}{2},3,\tfrac{7}{2}\right\},\left\{\tfrac{3}{2},\tfrac{7}{2},10\right\},\left\{\tfrac{5}{2},\tfrac{7}{2},5\right\},\left\{\tfrac{7}{2},5,6\right\},\left\{3,\tfrac{7}{2},5\right\},\left\{\tfrac{7}{2},5,10\right\}$
$\left\{\tfrac{5}{2},\tfrac{7}{2},12\right\},\left\{\tfrac{7}{2},6,12\right\},\left\{3,\tfrac{7}{2},12\right\},\left\{\tfrac{7}{2},10,12\right\},\left\{\tfrac{5}{2},\tfrac{7}{2},4\right\},\left\{\tfrac{7}{2},4,6\right\},\left\{3,\tfrac{7}{2},4\right\},\left\{\tfrac{7}{2},4,10\right\}$
$\tfrac{18}{5}$$\left\{\tfrac{6}{5},\tfrac{12}{5},\tfrac{18}{5}\right\},\left\{\tfrac{6}{5},\tfrac{18}{5},6\right\},\left\{\tfrac{12}{5},\tfrac{18}{5},12\right\},\left\{\tfrac{18}{5},6,12\right\},\left\{\tfrac{12}{5},\tfrac{18}{5},8\right\},\left\{\tfrac{18}{5},6,8\right\}$
$\tfrac{9}{2}$$\left\{\tfrac{3}{2},3,\tfrac{9}{2}\right\},\left\{\tfrac{3}{2},\tfrac{9}{2},12\right\},\left\{\tfrac{3}{2},\tfrac{9}{2},8\right\},\left\{3,\tfrac{9}{2},6\right\},\left\{\tfrac{9}{2},6,12\right\},\left\{\tfrac{9}{2},6,8\right\},\left\{3,\tfrac{9}{2},5\right\},\left\{\tfrac{9}{2},5,12\right\},\left\{\tfrac{9}{2},5,8\right\}$
$\tfrac{14}{3}$$\left\{\tfrac{4}{3},\tfrac{8}{3},\tfrac{14}{3}\right\},\left\{\tfrac{4}{3},\tfrac{14}{3},12\right\},\left\{\tfrac{4}{3},4,\tfrac{14}{3}\right\},\left\{\tfrac{8}{3},\tfrac{14}{3},6\right\},\left\{\tfrac{14}{3},6,12\right\},\left\{4,\tfrac{14}{3},6\right\},\left\{\tfrac{8}{3},\tfrac{14}{3},10\right\},\left\{\tfrac{14}{3},10,12\right\},\left\{4,\tfrac{14}{3},10\right\}$
$7$$\left\{3,4,7\right\},\left\{3,5,7\right\},\left\{3,7,12\right\},\left\{4,7,10\right\},\left\{5,7,10\right\},\left\{7,10,12\right\},\left\{4,6,7\right\},\left\{5,6,7\right\},\left\{6,7,12\right\}$
$9$$\left\{3,5,9\right\},\left\{3,6,9\right\},\left\{5,9,12\right\},\left\{6,9,12\right\},\left\{5,8,9\right\},\left\{6,8,9\right\}$
$14$$\left\{4,6,14\right\},\left\{4,10,14\right\},\left\{6,12,14\right\},\left\{10,12,14\right\}$
$18$$\left\{6,8,18\right\},\left\{6,12,18\right\}$
Table 4: The genuinely rank-$3$ sets of scaling dimensions with $\tfrac{7}{4}\leq\Delta_{\mathrm{new}}\leq 18$. The two triples in blue are repeated from table 3 since $7/5$ is also a genuinely rank-3 dimension.
$\Delta_{\mathrm{new}}$$\{\Delta_{1},\Delta_{2},\Delta_{3},\Delta_{4}\}$
$\tfrac{30}{29}$$\left\{\tfrac{30}{29},8,12,18\right\},\left\{\tfrac{30}{29},8,12,14\right\}$
$\tfrac{24}{23}$$\left\{\tfrac{24}{23},6,14,18\right\},\left\{\tfrac{24}{23},6,12,18\right\},\left\{\tfrac{24}{23},6,8,14\right\},\left\{\tfrac{24}{23},6,8,12\right\}$
$\tfrac{20}{19}$$\left\{\tfrac{20}{19},12,14,18\right\},\left\{\tfrac{20}{19},10,14,18\right\},\left\{\tfrac{20}{19},8,12,18\right\},\left\{\tfrac{20}{19},\tfrac{30}{19},8,18\right\},\left\{\tfrac{20}{19},8,10,18\right\}$
$\left\{\tfrac{20}{19},4,12,14\right\},\left\{\tfrac{20}{19},4,10,14\right\},\left\{\tfrac{20}{19},4,8,12\right\},\left\{\tfrac{20}{19},4,8,10\right\}$
$\tfrac{16}{15}$$\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{4}{3},\tfrac{18}{5}\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{4}{3},\tfrac{14}{3}\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{4}{3},10\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{4}{3},\tfrac{8}{5}\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{4}{3},\tfrac{8}{3}\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{4}{3},8\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{12}{5},\tfrac{18}{5}\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{12}{5},\tfrac{14}{3}\right\}$
$\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{12}{5},10\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{8}{5},\tfrac{12}{5}\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{12}{5},\tfrac{8}{3}\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{12}{5},8\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{18}{5},12\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{14}{3},12\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},10,12\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{8}{5},12\right\}$
$\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{8}{3},12\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},8,12\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{14}{5},\tfrac{18}{5}\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{18}{5},6\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{14}{5},\tfrac{14}{3}\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{14}{3},6\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{14}{5},10\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},6,10\right\}$
$\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{8}{5},\tfrac{14}{5}\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{8}{5},6\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{8}{3},\tfrac{14}{5}\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{8}{3},6\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},\tfrac{14}{5},8\right\},\left\{\tfrac{16}{15},\tfrac{6}{5},6,8\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{10}{3},\tfrac{18}{5}\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{10}{3},\tfrac{14}{3}\right\}$
$\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{10}{3},10\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{8}{5},\tfrac{10}{3}\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{8}{3},\tfrac{10}{3}\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{10}{3},8\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},\tfrac{10}{3},\tfrac{18}{5}\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},\tfrac{10}{3},\tfrac{14}{3}\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},\tfrac{10}{3},10\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},\tfrac{12}{5},\tfrac{10}{3}\right\}$
$\left\{\tfrac{16}{15},\tfrac{12}{5},\tfrac{8}{3},\tfrac{10}{3}\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},\tfrac{10}{3},8\right\},\left\{\tfrac{16}{15},\tfrac{10}{3},\tfrac{18}{5},12\right\},\left\{\tfrac{16}{15},\tfrac{10}{3},\tfrac{14}{3},12\right\},\left\{\tfrac{16}{15},\tfrac{10}{3},10,12\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},\tfrac{10}{3},12\right\},\left\{\tfrac{16}{15},\tfrac{8}{3},\tfrac{10}{3},12\right\},\left\{\tfrac{16}{15},\tfrac{10}{3},8,12\right\}$
$\left\{\tfrac{16}{15},\tfrac{14}{5},\tfrac{10}{3},\tfrac{18}{5}\right\},\left\{\tfrac{16}{15},\tfrac{10}{3},\tfrac{18}{5},6\right\},\left\{\tfrac{16}{15},\tfrac{14}{5},\tfrac{10}{3},\tfrac{14}{3}\right\},\left\{\tfrac{16}{15},\tfrac{10}{3},\tfrac{14}{3},6\right\},\left\{\tfrac{16}{15},\tfrac{14}{5},\tfrac{10}{3},10\right\},\left\{\tfrac{16}{15},\tfrac{10}{3},6,10\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},\tfrac{14}{5},\tfrac{10}{3}\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},\tfrac{10}{3},6\right\}$
$\left\{\tfrac{16}{15},\tfrac{8}{3},\tfrac{14}{5},\tfrac{10}{3}\right\},\left\{\tfrac{16}{15},\tfrac{8}{3},\tfrac{10}{3},6\right\},\left\{\tfrac{16}{15},\tfrac{14}{5},\tfrac{10}{3},8\right\},\left\{\tfrac{16}{15},\tfrac{10}{3},6,8\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{18}{5},14\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{14}{3},14\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},10,14\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{8}{5},14\right\}$
$\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{8}{3},14\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},8,14\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},\tfrac{18}{5},14\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},\tfrac{14}{3},14\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},10,14\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},\tfrac{12}{5},14\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},\tfrac{8}{3},14\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},8,14\right\}$
$\left\{\tfrac{16}{15},\tfrac{18}{5},12,14\right\},\left\{\tfrac{16}{15},\tfrac{14}{3},12,14\right\},\left\{\tfrac{16}{15},10,12,14\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},12,14\right\},\left\{\tfrac{16}{15},\tfrac{8}{3},12,14\right\},\left\{\tfrac{16}{15},8,12,14\right\},\left\{\tfrac{16}{15},\tfrac{14}{5},\tfrac{18}{5},14\right\},\left\{\tfrac{16}{15},\tfrac{18}{5},6,14\right\}$
$\left\{\tfrac{16}{15},\tfrac{14}{5},\tfrac{14}{3},14\right\},\left\{\tfrac{16}{15},\tfrac{14}{3},6,14\right\},\left\{\tfrac{16}{15},\tfrac{14}{5},10,14\right\},\left\{\tfrac{16}{15},6,10,14\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},\tfrac{14}{5},14\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},6,14\right\},\left\{\tfrac{16}{15},\tfrac{8}{3},\tfrac{14}{5},14\right\},\left\{\tfrac{16}{15},\tfrac{8}{3},6,14\right\}$
$\left\{\tfrac{16}{15},\tfrac{14}{5},8,14\right\},\left\{\tfrac{16}{15},6,8,14\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{18}{5},4\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},4,\tfrac{14}{3}\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},4,10\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{8}{5},4\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},\tfrac{8}{3},4\right\},\left\{\tfrac{16}{15},\tfrac{4}{3},4,8\right\}$
$\left\{\tfrac{16}{15},\tfrac{12}{5},\tfrac{18}{5},4\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},4,\tfrac{14}{3}\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},4,10\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},\tfrac{12}{5},4\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},\tfrac{8}{3},4\right\},\left\{\tfrac{16}{15},\tfrac{12}{5},4,8\right\},\left\{\tfrac{16}{15},\tfrac{18}{5},4,12\right\},\left\{\tfrac{16}{15},4,\tfrac{14}{3},12\right\}$
$\left\{\tfrac{16}{15},4,10,12\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},4,12\right\},\left\{\tfrac{16}{15},\tfrac{8}{3},4,12\right\},\left\{\tfrac{16}{15},4,8,12\right\},\left\{\tfrac{16}{15},\tfrac{14}{5},\tfrac{18}{5},4\right\},\left\{\tfrac{16}{15},\tfrac{18}{5},4,6\right\},\left\{\tfrac{16}{15},\tfrac{14}{5},4,\tfrac{14}{3}\right\},\left\{\tfrac{16}{15},4,\tfrac{14}{3},6\right\}$
$\left\{\tfrac{16}{15},\tfrac{14}{5},4,10\right\},\left\{\tfrac{16}{15},4,6,10\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},\tfrac{14}{5},4\right\},\left\{\tfrac{16}{15},\tfrac{8}{5},4,6\right\},\left\{\tfrac{16}{15},\tfrac{8}{3},\tfrac{14}{5},4\right\},\left\{\tfrac{16}{15},\tfrac{8}{3},4,6\right\},\left\{\tfrac{16}{15},\tfrac{14}{5},4,8\right\},\left\{\tfrac{16}{15},4,6,8\right\}$
Table 5: The genuinely rank-$4$ sets of scaling dimensions with $\tfrac{30}{29}\leq\Delta_{\mathrm{new}}\leq\tfrac{16}{15}$.
$\Delta_{\mathrm{new}}$$\{\Delta_{1},\Delta_{2},\Delta_{3},\Delta_{4}\}$
$\tfrac{15}{14}$$\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{9}{7},\tfrac{3}{2}\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{9}{7},\tfrac{18}{7}\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{9}{7},9\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{9}{7},8\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{3}{2},\tfrac{9}{2}\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{18}{7},\tfrac{9}{2}\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{9}{2},9\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{9}{2},8\right\}$
$\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{3}{2},12\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{18}{7},12\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},9,12\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},8,12\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{3}{2},5\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},\tfrac{18}{7},5\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},5,9\right\},\left\{\tfrac{15}{14},\tfrac{8}{7},5,8\right\}$
$\left\{\tfrac{15}{14},\tfrac{9}{7},\tfrac{3}{2},14\right\},\left\{\tfrac{15}{14},\tfrac{9}{7},\tfrac{18}{7},14\right\},\left\{\tfrac{15}{14},\tfrac{9}{7},9,14\right\},\left\{\tfrac{15}{14},\tfrac{9}{7},8,14\right\},\left\{\tfrac{15}{14},\tfrac{3}{2},\tfrac{9}{2},14\right\},\left\{\tfrac{15}{14},\tfrac{18}{7},\tfrac{9}{2},14\right\},\left\{\tfrac{15}{14},\tfrac{9}{2},9,14\right\},\left\{\tfrac{15}{14},\tfrac{9}{2},8,14\right\}$
$\left\{\tfrac{15}{14},\tfrac{3}{2},12,14\right\},\left\{\tfrac{15}{14},\tfrac{18}{7},12,14\right\},\left\{\tfrac{15}{14},9,12,14\right\},\left\{\tfrac{15}{14},8,12,14\right\},\left\{\tfrac{15}{14},\tfrac{3}{2},5,14\right\},\left\{\tfrac{15}{14},\tfrac{18}{7},5,14\right\},\left\{\tfrac{15}{14},5,9,14\right\},\left\{\tfrac{15}{14},5,8,14\right\}$
$\left\{\tfrac{15}{14},\tfrac{9}{7},\tfrac{3}{2},3\right\},\left\{\tfrac{15}{14},\tfrac{9}{7},\tfrac{3}{2},18\right\},\left\{\tfrac{15}{14},\tfrac{9}{7},\tfrac{18}{7},3\right\},\left\{\tfrac{15}{14},\tfrac{9}{7},\tfrac{18}{7},18\right\},\left\{\tfrac{15}{14},\tfrac{9}{7},3,9\right\},\left\{\tfrac{15}{14},\tfrac{9}{7},9,18\right\},\left\{\tfrac{15}{14},\tfrac{9}{7},3,8\right\},\left\{\tfrac{15}{14},\tfrac{9}{7},8,18\right\}$
$\left\{\tfrac{15}{14},\tfrac{3}{2},3,\tfrac{9}{2}\right\},\left\{\tfrac{15}{14},\tfrac{3}{2},\tfrac{9}{2},18\right\},\left\{\tfrac{15}{14},\tfrac{18}{7},3,\tfrac{9}{2}\right\},\left\{\tfrac{15}{14},\tfrac{18}{7},\tfrac{9}{2},18\right\},\left\{\tfrac{15}{14},3,\tfrac{9}{2},9\right\},\left\{\tfrac{15}{14},\tfrac{9}{2},9,18\right\},\left\{\tfrac{15}{14},3,\tfrac{9}{2},8\right\},\left\{\tfrac{15}{14},\tfrac{9}{2},8,18\right\}$
$\left\{\tfrac{15}{14},\tfrac{3}{2},3,12\right\},\left\{\tfrac{15}{14},\tfrac{3}{2},12,18\right\},\left\{\tfrac{15}{14},\tfrac{18}{7},3,12\right\},\left\{\tfrac{15}{14},\tfrac{18}{7},12,18\right\},\left\{\tfrac{15}{14},3,9,12\right\},\left\{\tfrac{15}{14},9,12,18\right\},\left\{\tfrac{15}{14},3,8,12\right\},\left\{\tfrac{15}{14},8,12,18\right\}$
$\left\{\tfrac{15}{14},\tfrac{3}{2},3,5\right\},\left\{\tfrac{15}{14},\tfrac{3}{2},5,18\right\},\left\{\tfrac{15}{14},\tfrac{18}{7},3,5\right\},\left\{\tfrac{15}{14},\tfrac{18}{7},5,18\right\},\left\{\tfrac{15}{14},3,5,9\right\},\left\{\tfrac{15}{14},5,9,18\right\},\left\{\tfrac{15}{14},3,5,8\right\},\left\{\tfrac{15}{14},5,8,18\right\}$
$\tfrac{15}{13}$$\left\{\tfrac{14}{13},\tfrac{15}{13},5,14\right\},\left\{\tfrac{14}{13},\tfrac{15}{13},12,14\right\},\left\{\tfrac{14}{13},\tfrac{15}{13},3,5\right\},\left\{\tfrac{14}{13},\tfrac{15}{13},5,18\right\},\left\{\tfrac{14}{13},\tfrac{15}{13},3,12\right\},\left\{\tfrac{14}{13},\tfrac{15}{13},12,18\right\}$
$\left\{\tfrac{15}{13},5,8,14\right\},\left\{\tfrac{15}{13},8,12,14\right\},\left\{\tfrac{15}{13},3,5,8\right\},\left\{\tfrac{15}{13},5,8,18\right\},\left\{\tfrac{15}{13},3,8,12\right\},\left\{\tfrac{15}{13},8,12,18\right\}$
$\left\{\tfrac{15}{13},5,9,14\right\},\left\{\tfrac{15}{13},9,12,14\right\},\left\{\tfrac{15}{13},3,5,9\right\},\left\{\tfrac{15}{13},5,9,18\right\},\left\{\tfrac{15}{13},3,9,12\right\},\left\{\tfrac{15}{13},9,12,18\right\}$
$\tfrac{20}{17}$$\left\{\tfrac{18}{17},\tfrac{20}{17},12,18\right\},\left\{\tfrac{18}{17},\tfrac{20}{17},4,12\right\},\left\{\tfrac{18}{17},\tfrac{20}{17},10,18\right\},\left\{\tfrac{18}{17},\tfrac{20}{17},4,10\right\},\left\{\tfrac{20}{17},12,14,18\right\},\left\{\tfrac{20}{17},4,12,14\right\}$
$\left\{\tfrac{20}{17},10,14,18\right\},\left\{\tfrac{20}{17},4,10,14\right\},\left\{\tfrac{20}{17},8,12,18\right\},\left\{\tfrac{20}{17},4,8,12\right\},\left\{\tfrac{20}{17},8,10,18\right\},\left\{\tfrac{20}{17},4,8,10\right\}$
$\tfrac{16}{13}$$\left\{\tfrac{14}{13},\tfrac{16}{13},\tfrac{18}{13},4\right\},\left\{\tfrac{14}{13},\tfrac{16}{13},\tfrac{18}{13},14\right\},\left\{\tfrac{14}{13},\tfrac{16}{13},4,10\right\},\left\{\tfrac{14}{13},\tfrac{16}{13},10,14\right\},\left\{\tfrac{14}{13},\tfrac{16}{13},\tfrac{20}{13},8\right\},\left\{\tfrac{14}{13},\tfrac{16}{13},4,8\right\},\left\{\tfrac{14}{13},\tfrac{16}{13},8,14\right\}$
$\left\{\tfrac{16}{13},\tfrac{18}{13},4,6\right\},\left\{\tfrac{16}{13},\tfrac{18}{13},6,14\right\},\left\{\tfrac{16}{13},4,6,10\right\},\left\{\tfrac{16}{13},6,10,14\right\},\left\{\tfrac{16}{13},4,6,8\right\},\left\{\tfrac{16}{13},6,8,14\right\}$
$\left\{\tfrac{16}{13},\tfrac{18}{13},4,12\right\},\left\{\tfrac{16}{13},\tfrac{18}{13},12,14\right\},\left\{\tfrac{16}{13},4,10,12\right\},\left\{\tfrac{16}{13},10,12,14\right\},\left\{\tfrac{16}{13},4,8,12\right\},\left\{\tfrac{16}{13},8,12,14\right\}$
$\tfrac{24}{19}$$\left\{\tfrac{24}{19},6,14,18\right\},\left\{\tfrac{24}{19},6,8,14\right\},\left\{\tfrac{24}{19},6,12,18\right\},\left\{\tfrac{24}{19},6,8,12\right\}$
Table 6: The genuinely rank-$4$ sets of scaling dimensions with $\tfrac{15}{14}\leq\Delta_{\mathrm{new}}\leq\tfrac{16}{13}$.
$\Delta_{\mathrm{new}}$$\{\Delta_{1},\Delta_{2},\Delta_{3},\Delta_{4}\}$
$\tfrac{30}{23}$$\left\{\tfrac{30}{23},8,12,18\right\},\left\{\tfrac{30}{23},8,12,14\right\}$
$\tfrac{15}{11}$$\left\{\tfrac{12}{11},\tfrac{15}{11},\tfrac{18}{11},8\right\},\left\{\tfrac{12}{11},\tfrac{15}{11},3,8\right\},\left\{\tfrac{12}{11},\tfrac{15}{11},8,18\right\},\left\{\tfrac{12}{11},\tfrac{15}{11},8,14\right\},\left\{\tfrac{12}{11},\tfrac{15}{11},\tfrac{18}{11},9\right\},\left\{\tfrac{12}{11},\tfrac{15}{11},3,9\right\},\left\{\tfrac{12}{11},\tfrac{15}{11},9,18\right\},\left\{\tfrac{12}{11},\tfrac{15}{11},9,14\right\}$
$\left\{\tfrac{15}{11},\tfrac{18}{11},8,12\right\},\left\{\tfrac{15}{11},3,8,12\right\},\left\{\tfrac{15}{11},8,12,18\right\},\left\{\tfrac{15}{11},8,12,14\right\},\left\{\tfrac{15}{11},\tfrac{18}{11},9,12\right\},\left\{\tfrac{15}{11},3,9,12\right\},\left\{\tfrac{15}{11},9,12,18\right\},\left\{\tfrac{15}{11},9,12,14\right\}$
$\left\{\tfrac{15}{11},\tfrac{18}{11},5,8\right\},\left\{\tfrac{15}{11},3,5,8\right\},\left\{\tfrac{15}{11},5,8,18\right\},\left\{\tfrac{15}{11},5,8,14\right\},\left\{\tfrac{15}{11},\tfrac{18}{11},5,9\right\},\left\{\tfrac{15}{11},3,5,9\right\},\left\{\tfrac{15}{11},5,9,18\right\},\left\{\tfrac{15}{11},5,9,14\right\}$
$\tfrac{24}{17}$$\left\{\tfrac{18}{17},\tfrac{24}{17},\tfrac{30}{17},14\right\},\left\{\tfrac{18}{17},\tfrac{24}{17},6,14\right\},\left\{\tfrac{18}{17},\tfrac{24}{17},6,12\right\},\left\{\tfrac{24}{17},6,14,18\right\},\left\{\tfrac{24}{17},6,12,18\right\},\left\{\tfrac{24}{17},\tfrac{30}{17},8,14\right\},\left\{\tfrac{24}{17},6,8,14\right\},\left\{\tfrac{24}{17},6,8,12\right\}$
$\tfrac{16}{11}$$\left\{\tfrac{12}{11},\tfrac{14}{11},\tfrac{16}{11},\tfrac{18}{11}\right\},\left\{\tfrac{12}{11},\tfrac{14}{11},\tfrac{16}{11},6\right\},\left\{\tfrac{12}{11},\tfrac{14}{11},\tfrac{16}{11},12\right\},\left\{\tfrac{12}{11},\tfrac{16}{11},\tfrac{18}{11},10\right\},\left\{\tfrac{12}{11},\tfrac{16}{11},6,10\right\},\left\{\tfrac{12}{11},\tfrac{16}{11},10,12\right\},\left\{\tfrac{12}{11},\tfrac{16}{11},\tfrac{18}{11},\tfrac{24}{11}\right\},\left\{\tfrac{12}{11},\tfrac{16}{11},\tfrac{18}{11},8\right\}$
$\left\{\tfrac{12}{11},\tfrac{16}{11},\tfrac{24}{11},6\right\},\left\{\tfrac{12}{11},\tfrac{16}{11},6,8\right\},\left\{\tfrac{12}{11},\tfrac{16}{11},8,12\right\},\left\{\tfrac{14}{11},\tfrac{16}{11},\tfrac{18}{11},4\right\},\left\{\tfrac{14}{11},\tfrac{16}{11},4,6\right\},\left\{\tfrac{14}{11},\tfrac{16}{11},4,12\right\},\left\{\tfrac{16}{11},\tfrac{18}{11},4,10\right\},\left\{\tfrac{16}{11},4,6,10\right\}$
$\left\{\tfrac{16}{11},4,10,12\right\},\left\{\tfrac{16}{11},\tfrac{18}{11},4,8\right\},\left\{\tfrac{16}{11},4,6,8\right\},\left\{\tfrac{16}{11},4,8,12\right\},\left\{\tfrac{14}{11},\tfrac{16}{11},\tfrac{18}{11},14\right\},\left\{\tfrac{14}{11},\tfrac{16}{11},6,14\right\},\left\{\tfrac{14}{11},\tfrac{16}{11},12,14\right\},\left\{\tfrac{16}{11},\tfrac{18}{11},10,14\right\}$
$\left\{\tfrac{16}{11},6,10,14\right\},\left\{\tfrac{16}{11},10,12,14\right\},\left\{\tfrac{16}{11},\tfrac{18}{11},\tfrac{24}{11},14\right\},\left\{\tfrac{16}{11},\tfrac{18}{11},8,14\right\},\left\{\tfrac{16}{11},\tfrac{24}{11},6,14\right\},\left\{\tfrac{16}{11},6,8,14\right\},\left\{\tfrac{16}{11},8,12,14\right\}$
$\tfrac{20}{13}$$\left\{\tfrac{14}{13},\tfrac{16}{13},\tfrac{20}{13},8\right\},\left\{\tfrac{14}{13},\tfrac{20}{13},12,14\right\},\left\{\tfrac{14}{13},\tfrac{20}{13},8,12\right\},\left\{\tfrac{14}{13},\tfrac{20}{13},\tfrac{30}{13},14\right\},\left\{\tfrac{14}{13},\tfrac{20}{13},10,14\right\},\left\{\tfrac{14}{13},\tfrac{20}{13},8,10\right\},\left\{\tfrac{20}{13},12,14,18\right\},\left\{\tfrac{20}{13},\tfrac{24}{13},12,18\right\}$
$\left\{\tfrac{20}{13},8,12,18\right\},\left\{\tfrac{20}{13},10,14,18\right\},\left\{\tfrac{20}{13},\tfrac{30}{13},8,18\right\},\left\{\tfrac{20}{13},8,10,18\right\},\left\{\tfrac{20}{13},4,12,14\right\},\left\{\tfrac{20}{13},4,8,12\right\},\left\{\tfrac{20}{13},4,10,14\right\},\left\{\tfrac{20}{13},4,8,10\right\}$
$\tfrac{30}{19}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{20}{19},\tfrac{30}{19},8,18\right\}}},\left\{\tfrac{30}{19},8,12,14\right\},\left\{\tfrac{30}{19},8,12,18\right\}$
$\tfrac{30}{17}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{18}{17},\tfrac{24}{17},\tfrac{30}{17},14\right\}}},\left\{\tfrac{18}{17},\tfrac{30}{17},12,14\right\},\left\{\tfrac{18}{17},\tfrac{30}{17},12,18\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{24}{17},\tfrac{30}{17},8,14\right\}}},\left\{\tfrac{30}{17},8,12,14\right\},\left\{\tfrac{30}{17},8,12,18\right\}$
Table 7: The genuinely rank-$4$ sets of scaling dimensions with $\tfrac{30}{23}\leq\Delta_{\mathrm{new}}\leq\tfrac{30}{17}$.
$\Delta_{\mathrm{new}}$$\{\Delta_{1},\Delta_{2},\Delta_{3},\Delta_{4}\}$
$\tfrac{16}{9}$$\left\{\tfrac{10}{9},\tfrac{4}{3},\tfrac{14}{9},\tfrac{16}{9}\right\},\left\{\tfrac{10}{9},\tfrac{4}{3},\tfrac{16}{9},\tfrac{10}{3}\right\},\left\{\tfrac{10}{9},\tfrac{4}{3},\tfrac{16}{9},14\right\},\left\{\tfrac{10}{9},\tfrac{4}{3},\tfrac{16}{9},\tfrac{20}{9}\right\},\left\{\tfrac{10}{9},\tfrac{4}{3},\tfrac{16}{9},4\right\},\left\{\tfrac{10}{9},\tfrac{14}{9},\tfrac{16}{9},12\right\},\left\{\tfrac{10}{9},\tfrac{16}{9},\tfrac{10}{3},12\right\},\left\{\tfrac{10}{9},\tfrac{16}{9},12,14\right\}$
$\left\{\tfrac{10}{9},\tfrac{16}{9},4,12\right\},\left\{\tfrac{10}{9},\tfrac{14}{9},\tfrac{16}{9},6\right\},\left\{\tfrac{10}{9},\tfrac{16}{9},\tfrac{10}{3},6\right\},\left\{\tfrac{10}{9},\tfrac{16}{9},6,14\right\},\left\{\tfrac{10}{9},\tfrac{16}{9},4,6\right\},\left\{\tfrac{4}{3},\tfrac{14}{9},\tfrac{16}{9},\tfrac{14}{3}\right\},\left\{\tfrac{4}{3},\tfrac{16}{9},\tfrac{10}{3},\tfrac{14}{3}\right\},\left\{\tfrac{4}{3},\tfrac{16}{9},\tfrac{14}{3},14\right\}$
$\left\{\tfrac{4}{3},\tfrac{16}{9},4,\tfrac{14}{3}\right\},\left\{\tfrac{14}{9},\tfrac{16}{9},\tfrac{14}{3},12\right\},\left\{\tfrac{16}{9},\tfrac{10}{3},\tfrac{14}{3},12\right\},\left\{\tfrac{16}{9},\tfrac{14}{3},12,14\right\},\left\{\tfrac{16}{9},4,\tfrac{14}{3},12\right\},\left\{\tfrac{14}{9},\tfrac{16}{9},\tfrac{14}{3},6\right\},\left\{\tfrac{16}{9},\tfrac{10}{3},\tfrac{14}{3},6\right\},\left\{\tfrac{16}{9},\tfrac{14}{3},6,14\right\}$
$\left\{\tfrac{16}{9},4,\tfrac{14}{3},6\right\},\left\{\tfrac{4}{3},\tfrac{14}{9},\tfrac{16}{9},10\right\},\left\{\tfrac{4}{3},\tfrac{16}{9},\tfrac{10}{3},10\right\},\left\{\tfrac{4}{3},\tfrac{16}{9},10,14\right\},\left\{\tfrac{4}{3},\tfrac{16}{9},\tfrac{20}{9},10\right\},\left\{\tfrac{4}{3},\tfrac{16}{9},4,10\right\},\left\{\tfrac{14}{9},\tfrac{16}{9},10,12\right\},\left\{\tfrac{16}{9},\tfrac{10}{3},10,12\right\}$
$\left\{\tfrac{16}{9},10,12,14\right\},\left\{\tfrac{16}{9},4,10,12\right\},\left\{\tfrac{14}{9},\tfrac{16}{9},6,10\right\},\left\{\tfrac{16}{9},\tfrac{10}{3},6,10\right\},\left\{\tfrac{16}{9},6,10,14\right\},\left\{\tfrac{16}{9},4,6,10\right\},\left\{\tfrac{4}{3},\tfrac{14}{9},\tfrac{16}{9},\tfrac{8}{3}\right\},\left\{\tfrac{4}{3},\tfrac{14}{9},\tfrac{16}{9},8\right\}$
$\left\{\tfrac{4}{3},\tfrac{16}{9},\tfrac{8}{3},\tfrac{10}{3}\right\},\left\{\tfrac{4}{3},\tfrac{16}{9},\tfrac{10}{3},8\right\},\left\{\tfrac{4}{3},\tfrac{16}{9},\tfrac{8}{3},14\right\},\left\{\tfrac{4}{3},\tfrac{16}{9},8,14\right\},\left\{\tfrac{4}{3},\tfrac{16}{9},\tfrac{8}{3},4\right\},\left\{\tfrac{4}{3},\tfrac{16}{9},4,8\right\},\left\{\tfrac{14}{9},\tfrac{16}{9},\tfrac{8}{3},12\right\},\left\{\tfrac{14}{9},\tfrac{16}{9},8,12\right\}$
$\left\{\tfrac{16}{9},\tfrac{8}{3},\tfrac{10}{3},12\right\},\left\{\tfrac{16}{9},\tfrac{10}{3},8,12\right\},\left\{\tfrac{16}{9},\tfrac{8}{3},12,14\right\},\left\{\tfrac{16}{9},8,12,14\right\},\left\{\tfrac{16}{9},\tfrac{20}{9},8,12\right\},\left\{\tfrac{16}{9},\tfrac{8}{3},4,12\right\},\left\{\tfrac{16}{9},4,8,12\right\},\left\{\tfrac{14}{9},\tfrac{16}{9},\tfrac{8}{3},6\right\}$
$\left\{\tfrac{14}{9},\tfrac{16}{9},6,8\right\},\left\{\tfrac{16}{9},\tfrac{8}{3},\tfrac{10}{3},6\right\},\left\{\tfrac{16}{9},\tfrac{10}{3},6,8\right\},\left\{\tfrac{16}{9},\tfrac{8}{3},6,14\right\},\left\{\tfrac{16}{9},6,8,14\right\},\left\{\tfrac{16}{9},\tfrac{8}{3},4,6\right\},\left\{\tfrac{16}{9},4,6,8\right\}$
$\tfrac{20}{11}$$\left\{\tfrac{12}{11},\tfrac{14}{11},\tfrac{18}{11},\tfrac{20}{11}\right\},\left\{\tfrac{12}{11},\tfrac{14}{11},\tfrac{20}{11},18\right\},\left\{\tfrac{12}{11},\tfrac{14}{11},\tfrac{20}{11},4\right\},\left\{\tfrac{12}{11},\tfrac{18}{11},\tfrac{20}{11},14\right\},\left\{\tfrac{12}{11},\tfrac{20}{11},14,18\right\},\left\{\tfrac{12}{11},\tfrac{20}{11},4,14\right\},\left\{\tfrac{12}{11},\tfrac{18}{11},\tfrac{20}{11},8\right\}$
$\left\{\tfrac{12}{11},\tfrac{20}{11},8,18\right\},\left\{\tfrac{12}{11},\tfrac{20}{11},4,8\right\},\left\{\tfrac{14}{11},\tfrac{18}{11},\tfrac{20}{11},12\right\},\left\{\tfrac{14}{11},\tfrac{20}{11},12,18\right\},\left\{\tfrac{14}{11},\tfrac{20}{11},4,12\right\},\left\{\tfrac{18}{11},\tfrac{20}{11},12,14\right\},\left\{\tfrac{20}{11},12,14,18\right\}$
$\left\{\tfrac{20}{11},4,12,14\right\},\left\{\tfrac{18}{11},\tfrac{20}{11},8,12\right\},\left\{\tfrac{20}{11},8,12,18\right\},\left\{\tfrac{20}{11},4,8,12\right\},\left\{\tfrac{14}{11},\tfrac{18}{11},\tfrac{20}{11},10\right\},\left\{\tfrac{14}{11},\tfrac{20}{11},10,18\right\},\left\{\tfrac{14}{11},\tfrac{20}{11},4,10\right\}$
$\left\{\tfrac{18}{11},\tfrac{20}{11},10,14\right\},\left\{\tfrac{20}{11},10,14,18\right\},\left\{\tfrac{20}{11},4,10,14\right\},\left\{\tfrac{18}{11},\tfrac{20}{11},8,10\right\},\left\{\tfrac{20}{11},8,10,18\right\},\left\{\tfrac{20}{11},4,8,10\right\}$
$\tfrac{24}{13}$$\left\{\tfrac{14}{13},\tfrac{18}{13},\tfrac{24}{13},6\right\},\left\{\tfrac{14}{13},\tfrac{24}{13},6,18\right\},\left\{\tfrac{14}{13},\tfrac{24}{13},6,8\right\},\left\{\tfrac{18}{13},\tfrac{24}{13},6,14\right\},\left\{\tfrac{24}{13},6,14,18\right\},\left\{\tfrac{24}{13},6,8,14\right\}$
$\left\{\tfrac{18}{13},\tfrac{24}{13},6,12\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{20}{13},\tfrac{24}{13},12,18\right\}}},\left\{\tfrac{24}{13},6,12,18\right\},\left\{\tfrac{24}{13},\tfrac{30}{13},8,12\right\},\left\{\tfrac{24}{13},6,8,12\right\}$
$\tfrac{15}{8}$$\left\{\tfrac{9}{8},\tfrac{5}{4},\tfrac{3}{2},\tfrac{15}{8}\right\},\left\{\tfrac{9}{8},\tfrac{5}{4},\tfrac{15}{8},9\right\},\left\{\tfrac{9}{8},\tfrac{5}{4},\tfrac{15}{8},8\right\},\left\{\tfrac{9}{8},\tfrac{3}{2},\tfrac{15}{8},5\right\},\left\{\tfrac{9}{8},\tfrac{15}{8},5,9\right\},\left\{\tfrac{9}{8},\tfrac{15}{8},5,8\right\},\left\{\tfrac{9}{8},\tfrac{3}{2},\tfrac{15}{8},\tfrac{9}{2}\right\},\left\{\tfrac{9}{8},\tfrac{3}{2},\tfrac{15}{8},12\right\}$
$\left\{\tfrac{9}{8},\tfrac{15}{8},\tfrac{9}{2},9\right\},\left\{\tfrac{9}{8},\tfrac{15}{8},9,12\right\},\left\{\tfrac{9}{8},\tfrac{15}{8},\tfrac{9}{2},8\right\},\left\{\tfrac{9}{8},\tfrac{15}{8},8,12\right\},\left\{\tfrac{5}{4},\tfrac{3}{2},\tfrac{15}{8},3\right\},\left\{\tfrac{5}{4},\tfrac{15}{8},3,9\right\},\left\{\tfrac{5}{4},\tfrac{15}{8},3,8\right\},\left\{\tfrac{3}{2},\tfrac{15}{8},3,5\right\}$
$\left\{\tfrac{15}{8},3,5,9\right\},\left\{\tfrac{15}{8},3,5,8\right\},\left\{\tfrac{3}{2},\tfrac{15}{8},3,\tfrac{9}{2}\right\},\left\{\tfrac{3}{2},\tfrac{15}{8},3,12\right\},\left\{\tfrac{15}{8},3,\tfrac{9}{2},9\right\},\left\{\tfrac{15}{8},3,9,12\right\},\left\{\tfrac{15}{8},3,\tfrac{9}{2},8\right\},\left\{\tfrac{15}{8},3,8,12\right\}$
$\left\{\tfrac{5}{4},\tfrac{3}{2},\tfrac{15}{8},18\right\},\left\{\tfrac{5}{4},\tfrac{15}{8},9,18\right\},\left\{\tfrac{5}{4},\tfrac{15}{8},8,18\right\},\left\{\tfrac{3}{2},\tfrac{15}{8},5,18\right\},\left\{\tfrac{15}{8},5,9,18\right\},\left\{\tfrac{15}{8},5,8,18\right\},\left\{\tfrac{3}{2},\tfrac{15}{8},\tfrac{9}{2},18\right\},\left\{\tfrac{3}{2},\tfrac{15}{8},12,18\right\}$
$\left\{\tfrac{15}{8},\tfrac{9}{2},9,18\right\},\left\{\tfrac{15}{8},9,12,18\right\},\left\{\tfrac{15}{8},\tfrac{9}{2},8,18\right\},\left\{\tfrac{15}{8},8,12,18\right\},\left\{\tfrac{5}{4},\tfrac{3}{2},\tfrac{15}{8},14\right\},\left\{\tfrac{5}{4},\tfrac{15}{8},9,14\right\},\left\{\tfrac{5}{4},\tfrac{15}{8},8,14\right\},\left\{\tfrac{3}{2},\tfrac{15}{8},5,14\right\}$
$\left\{\tfrac{15}{8},5,9,14\right\},\left\{\tfrac{15}{8},5,8,14\right\},\left\{\tfrac{3}{2},\tfrac{15}{8},\tfrac{9}{2},14\right\},\left\{\tfrac{3}{2},\tfrac{15}{8},12,14\right\},\left\{\tfrac{15}{8},\tfrac{9}{2},9,14\right\},\left\{\tfrac{15}{8},9,12,14\right\},\left\{\tfrac{15}{8},\tfrac{9}{2},8,14\right\},\left\{\tfrac{15}{8},8,12,14\right\}$
Table 8: The genuinely rank-$4$ sets of scaling dimensions with $\tfrac{16}{9}\leq\Delta_{\mathrm{new}}\leq\tfrac{15}{8}$.
$\Delta_{\mathrm{new}}$$\{\Delta_{1},\Delta_{2},\Delta_{3},\Delta_{4}\}$
$\tfrac{15}{7}$$\left\{\tfrac{8}{7},\tfrac{9}{7},\tfrac{15}{7},8\right\},\left\{\tfrac{8}{7},\tfrac{9}{7},\tfrac{15}{7},\tfrac{18}{7}\right\},\left\{\tfrac{8}{7},\tfrac{9}{7},\tfrac{15}{7},9\right\},\left\{\tfrac{8}{7},\tfrac{15}{7},8,12\right\},\left\{\tfrac{8}{7},\tfrac{15}{7},\tfrac{18}{7},12\right\},\left\{\tfrac{8}{7},\tfrac{15}{7},9,12\right\},\left\{\tfrac{8}{7},\tfrac{15}{7},5,8\right\},\left\{\tfrac{8}{7},\tfrac{15}{7},\tfrac{18}{7},5\right\}$
$\left\{\tfrac{8}{7},\tfrac{15}{7},5,9\right\},\left\{\tfrac{9}{7},\tfrac{15}{7},8,14\right\},\left\{\tfrac{9}{7},\tfrac{15}{7},\tfrac{18}{7},14\right\},\left\{\tfrac{9}{7},\tfrac{15}{7},9,14\right\},\left\{\tfrac{15}{7},8,12,14\right\},\left\{\tfrac{15}{7},\tfrac{18}{7},12,14\right\},\left\{\tfrac{15}{7},9,12,14\right\}$
$\left\{\tfrac{15}{7},5,8,14\right\},\left\{\tfrac{15}{7},\tfrac{18}{7},5,14\right\},\left\{\tfrac{15}{7},5,9,14\right\},\left\{\tfrac{9}{7},\tfrac{15}{7},3,8\right\},\left\{\tfrac{9}{7},\tfrac{15}{7},8,18\right\},\left\{\tfrac{9}{7},\tfrac{15}{7},\tfrac{18}{7},3\right\},\left\{\tfrac{9}{7},\tfrac{15}{7},\tfrac{18}{7},18\right\}$
$\left\{\tfrac{9}{7},\tfrac{15}{7},3,9\right\},\left\{\tfrac{9}{7},\tfrac{15}{7},9,18\right\},\left\{\tfrac{15}{7},3,8,12\right\},\left\{\tfrac{15}{7},8,12,18\right\},\left\{\tfrac{15}{7},\tfrac{18}{7},3,12\right\},\left\{\tfrac{15}{7},\tfrac{18}{7},12,18\right\},\left\{\tfrac{15}{7},3,9,12\right\}$
$\left\{\tfrac{15}{7},9,12,18\right\},\left\{\tfrac{15}{7},3,5,8\right\},\left\{\tfrac{15}{7},5,8,18\right\},\left\{\tfrac{15}{7},\tfrac{18}{7},3,5\right\},\left\{\tfrac{15}{7},\tfrac{18}{7},5,18\right\},\left\{\tfrac{15}{7},3,5,9\right\},\left\{\tfrac{15}{7},5,9,18\right\}$
$\tfrac{24}{11}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{12}{11},\tfrac{16}{11},\tfrac{18}{11},\tfrac{24}{11}\right\}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{12}{11},\tfrac{16}{11},\tfrac{24}{11},6\right\}}},\left\{\tfrac{12}{11},\tfrac{18}{11},\tfrac{24}{11},8\right\},\left\{\tfrac{12}{11},\tfrac{24}{11},6,8\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{12}{11},\tfrac{18}{11},\tfrac{24}{11},\tfrac{30}{11}\right\}}},\left\{\tfrac{12}{11},\tfrac{18}{11},\tfrac{24}{11},18\right\}$
$\left\{\tfrac{12}{11},\tfrac{24}{11},6,18\right\},\left\{\tfrac{18}{11},\tfrac{24}{11},8,12\right\},\left\{\tfrac{24}{11},6,8,12\right\},\left\{\tfrac{18}{11},\tfrac{24}{11},\tfrac{30}{11},12\right\},\left\{\tfrac{18}{11},\tfrac{24}{11},12,18\right\},\left\{\tfrac{24}{11},6,12,18\right\}$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{16}{11},\tfrac{18}{11},\tfrac{24}{11},14\right\}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{16}{11},\tfrac{24}{11},6,14\right\}}},\left\{\tfrac{18}{11},\tfrac{24}{11},8,14\right\},\left\{\tfrac{24}{11},6,8,14\right\},\left\{\tfrac{18}{11},\tfrac{24}{11},14,18\right\},\left\{\tfrac{24}{11},6,14,18\right\}$
$\tfrac{20}{9}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{10}{9},\tfrac{4}{3},\tfrac{16}{9},\tfrac{20}{9}\right\}}},\left\{\tfrac{10}{9},\tfrac{4}{3},\tfrac{20}{9},4\right\},\left\{\tfrac{10}{9},\tfrac{4}{3},\tfrac{20}{9},\tfrac{14}{3}\right\},\left\{\tfrac{10}{9},\tfrac{4}{3},\tfrac{20}{9},18\right\},\left\{\tfrac{10}{9},\tfrac{20}{9},4,8\right\},\left\{\tfrac{10}{9},\tfrac{20}{9},\tfrac{14}{3},8\right\},\left\{\tfrac{10}{9},\tfrac{20}{9},8,18\right\},\left\{\tfrac{10}{9},\tfrac{20}{9},4,14\right\}$
$\left\{\tfrac{10}{9},\tfrac{20}{9},\tfrac{14}{3},14\right\},\left\{\tfrac{10}{9},\tfrac{20}{9},14,18\right\},\left\{\tfrac{4}{3},\tfrac{20}{9},\tfrac{10}{3},4\right\},\left\{\tfrac{4}{3},\tfrac{20}{9},\tfrac{10}{3},\tfrac{14}{3}\right\},\left\{\tfrac{4}{3},\tfrac{20}{9},\tfrac{10}{3},18\right\},\left\{\tfrac{20}{9},\tfrac{10}{3},4,8\right\},\left\{\tfrac{20}{9},\tfrac{10}{3},\tfrac{14}{3},8\right\},\left\{\tfrac{20}{9},\tfrac{10}{3},8,18\right\}$
$\left\{\tfrac{20}{9},\tfrac{10}{3},4,14\right\},\left\{\tfrac{20}{9},\tfrac{10}{3},\tfrac{14}{3},14\right\},\left\{\tfrac{20}{9},\tfrac{10}{3},14,18\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{4}{3},\tfrac{16}{9},\tfrac{20}{9},10\right\}}},\left\{\tfrac{4}{3},\tfrac{20}{9},4,10\right\},\left\{\tfrac{4}{3},\tfrac{20}{9},\tfrac{14}{3},10\right\},\left\{\tfrac{4}{3},\tfrac{20}{9},10,18\right\},\left\{\tfrac{20}{9},4,8,10\right\}$
$\left\{\tfrac{20}{9},\tfrac{14}{3},8,10\right\},\left\{\tfrac{20}{9},8,10,18\right\},\left\{\tfrac{20}{9},4,10,14\right\},\left\{\tfrac{20}{9},\tfrac{14}{3},10,14\right\},\left\{\tfrac{20}{9},10,14,18\right\},\left\{\tfrac{4}{3},\tfrac{20}{9},4,12\right\},\left\{\tfrac{4}{3},\tfrac{20}{9},\tfrac{14}{3},12\right\},\left\{\tfrac{4}{3},\tfrac{20}{9},12,18\right\}$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{16}{9},\tfrac{20}{9},8,12\right\}}},\left\{\tfrac{20}{9},4,8,12\right\},\left\{\tfrac{20}{9},\tfrac{14}{3},8,12\right\},\left\{\tfrac{20}{9},8,12,18\right\},\left\{\tfrac{20}{9},4,12,14\right\},\left\{\tfrac{20}{9},\tfrac{14}{3},12,14\right\},\left\{\tfrac{20}{9},12,14,18\right\}$
$\tfrac{16}{7}$$\left\{\tfrac{8}{7},\tfrac{10}{7},\tfrac{12}{7},\tfrac{16}{7}\right\},\left\{\tfrac{8}{7},\tfrac{10}{7},\tfrac{16}{7},4\right\},\left\{\tfrac{8}{7},\tfrac{10}{7},\tfrac{16}{7},\tfrac{18}{7}\right\},\left\{\tfrac{8}{7},\tfrac{10}{7},\tfrac{16}{7},14\right\},\left\{\tfrac{8}{7},\tfrac{12}{7},\tfrac{16}{7},6\right\},\left\{\tfrac{8}{7},\tfrac{16}{7},4,6\right\},\left\{\tfrac{8}{7},\tfrac{16}{7},\tfrac{18}{7},6\right\},\left\{\tfrac{8}{7},\tfrac{16}{7},6,14\right\}$
$\left\{\tfrac{8}{7},\tfrac{12}{7},\tfrac{16}{7},12\right\},\left\{\tfrac{8}{7},\tfrac{16}{7},4,12\right\},\left\{\tfrac{8}{7},\tfrac{16}{7},\tfrac{18}{7},12\right\},\left\{\tfrac{8}{7},\tfrac{16}{7},12,14\right\},\left\{\tfrac{10}{7},\tfrac{12}{7},\tfrac{16}{7},8\right\},\left\{\tfrac{10}{7},\tfrac{16}{7},4,8\right\},\left\{\tfrac{10}{7},\tfrac{16}{7},\tfrac{18}{7},8\right\},\left\{\tfrac{10}{7},\tfrac{16}{7},8,14\right\}$
$\left\{\tfrac{12}{7},\tfrac{16}{7},6,8\right\},\left\{\tfrac{16}{7},4,6,8\right\},\left\{\tfrac{16}{7},\tfrac{18}{7},6,8\right\},\left\{\tfrac{16}{7},6,8,14\right\},\left\{\tfrac{12}{7},\tfrac{16}{7},8,12\right\},\left\{\tfrac{16}{7},4,8,12\right\},\left\{\tfrac{16}{7},\tfrac{18}{7},8,12\right\},\left\{\tfrac{16}{7},8,12,14\right\}$
$\left\{\tfrac{10}{7},\tfrac{12}{7},\tfrac{16}{7},10\right\},\left\{\tfrac{10}{7},\tfrac{16}{7},4,10\right\},\left\{\tfrac{10}{7},\tfrac{16}{7},\tfrac{18}{7},10\right\},\left\{\tfrac{10}{7},\tfrac{16}{7},10,14\right\},\left\{\tfrac{12}{7},\tfrac{16}{7},6,10\right\},\left\{\tfrac{16}{7},4,6,10\right\},\left\{\tfrac{16}{7},\tfrac{18}{7},6,10\right\}$
$\left\{\tfrac{16}{7},6,10,14\right\},\left\{\tfrac{12}{7},\tfrac{16}{7},10,12\right\},\left\{\tfrac{16}{7},\tfrac{20}{7},4,10\right\},\left\{\tfrac{16}{7},4,10,12\right\},\left\{\tfrac{16}{7},\tfrac{18}{7},10,12\right\},\left\{\tfrac{16}{7},10,12,14\right\}$
Table 9: The genuinely rank-$4$ sets of scaling dimensions with $\tfrac{15}{7}\leq\Delta_{\mathrm{new}}\leq\tfrac{16}{7}$.
$\Delta_{\mathrm{new}}$$\{\Delta_{1},\Delta_{2},\Delta_{3},\Delta_{4}\}$
$\tfrac{30}{13}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{14}{13},\tfrac{20}{13},\tfrac{30}{13},14\right\}}},\left\{\tfrac{14}{13},\tfrac{30}{13},12,18\right\},\left\{\tfrac{14}{13},\tfrac{30}{13},12,14\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{20}{13},\tfrac{30}{13},8,18\right\}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{24}{13},\tfrac{30}{13},8,12\right\}}},\left\{\tfrac{30}{13},8,12,18\right\},\left\{\tfrac{30}{13},8,12,14\right\}$
$\tfrac{30}{11}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{12}{11},\tfrac{18}{11},\tfrac{24}{11},\tfrac{30}{11}\right\}}},\left\{\tfrac{12}{11},\tfrac{18}{11},\tfrac{30}{11},8\right\},\left\{\tfrac{12}{11},\tfrac{30}{11},8,18\right\},\left\{\tfrac{12}{11},\tfrac{30}{11},8,14\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{18}{11},\tfrac{24}{11},\tfrac{30}{11},12\right\}}},\left\{\tfrac{18}{11},\tfrac{30}{11},8,12\right\},\left\{\tfrac{30}{11},8,12,18\right\},\left\{\tfrac{30}{11},8,12,14\right\}$
$\tfrac{20}{7}$$\left\{\tfrac{8}{7},\tfrac{10}{7},\tfrac{20}{7},8\right\},\left\{\tfrac{8}{7},\tfrac{10}{7},\tfrac{18}{7},\tfrac{20}{7}\right\},\left\{\tfrac{8}{7},\tfrac{10}{7},\tfrac{20}{7},14\right\},\left\{\tfrac{8}{7},\tfrac{20}{7},\tfrac{30}{7},8\right\},\left\{\tfrac{8}{7},\tfrac{18}{7},\tfrac{20}{7},\tfrac{30}{7}\right\},\left\{\tfrac{8}{7},\tfrac{20}{7},8,10\right\},\left\{\tfrac{8}{7},\tfrac{18}{7},\tfrac{20}{7},10\right\}$
$\left\{\tfrac{8}{7},\tfrac{20}{7},10,14\right\},\left\{\tfrac{8}{7},\tfrac{20}{7},8,12\right\},\left\{\tfrac{8}{7},\tfrac{18}{7},\tfrac{20}{7},12\right\},\left\{\tfrac{8}{7},\tfrac{20}{7},\tfrac{24}{7},14\right\},\left\{\tfrac{8}{7},\tfrac{20}{7},12,14\right\},\left\{\tfrac{10}{7},\tfrac{20}{7},4,8\right\},\left\{\tfrac{10}{7},\tfrac{18}{7},\tfrac{20}{7},4\right\}$
$\left\{\tfrac{10}{7},\tfrac{20}{7},4,14\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{16}{7},\tfrac{20}{7},4,10\right\}}},\left\{\tfrac{20}{7},4,8,10\right\},\left\{\tfrac{18}{7},\tfrac{20}{7},4,10\right\},\left\{\tfrac{20}{7},4,10,14\right\},\left\{\tfrac{20}{7},4,8,12\right\},\left\{\tfrac{18}{7},\tfrac{20}{7},4,12\right\}$
$\left\{\tfrac{20}{7},4,12,14\right\},\left\{\tfrac{10}{7},\tfrac{20}{7},8,18\right\},\left\{\tfrac{10}{7},\tfrac{18}{7},\tfrac{20}{7},18\right\},\left\{\tfrac{10}{7},\tfrac{20}{7},14,18\right\},\left\{\tfrac{20}{7},\tfrac{30}{7},8,18\right\},\left\{\tfrac{18}{7},\tfrac{20}{7},\tfrac{30}{7},18\right\},\left\{\tfrac{20}{7},8,10,18\right\}$
$\left\{\tfrac{18}{7},\tfrac{20}{7},10,18\right\},\left\{\tfrac{20}{7},10,14,18\right\},\left\{\tfrac{20}{7},8,12,18\right\},\left\{\tfrac{18}{7},\tfrac{20}{7},12,18\right\},\left\{\tfrac{20}{7},\tfrac{24}{7},14,18\right\},\left\{\tfrac{20}{7},12,14,18\right\}$
$\tfrac{16}{5}$$\left\{\tfrac{6}{5},\tfrac{8}{5},\tfrac{12}{5},\tfrac{16}{5}\right\},\left\{\tfrac{6}{5},\tfrac{8}{5},\tfrac{16}{5},12\right\},\left\{\tfrac{6}{5},\tfrac{8}{5},\tfrac{14}{5},\tfrac{16}{5}\right\},\left\{\tfrac{6}{5},\tfrac{8}{5},\tfrac{16}{5},6\right\},\left\{\tfrac{6}{5},\tfrac{12}{5},\tfrac{16}{5},\tfrac{24}{5}\right\},\left\{\tfrac{6}{5},\tfrac{16}{5},\tfrac{24}{5},12\right\},\left\{\tfrac{6}{5},\tfrac{12}{5},\tfrac{16}{5},8\right\},\left\{\tfrac{6}{5},\tfrac{16}{5},8,12\right\}$
$\left\{\tfrac{6}{5},\tfrac{14}{5},\tfrac{16}{5},8\right\},\left\{\tfrac{6}{5},\tfrac{16}{5},6,8\right\},\left\{\tfrac{6}{5},\tfrac{12}{5},\tfrac{16}{5},\tfrac{18}{5}\right\},\left\{\tfrac{6}{5},\tfrac{12}{5},\tfrac{16}{5},10\right\},\left\{\tfrac{6}{5},\tfrac{16}{5},\tfrac{18}{5},12\right\},\left\{\tfrac{6}{5},\tfrac{16}{5},10,12\right\},\left\{\tfrac{6}{5},\tfrac{14}{5},\tfrac{16}{5},\tfrac{18}{5}\right\},\left\{\tfrac{6}{5},\tfrac{14}{5},\tfrac{16}{5},10\right\}$
$\left\{\tfrac{6}{5},\tfrac{16}{5},\tfrac{18}{5},6\right\},\left\{\tfrac{6}{5},\tfrac{16}{5},6,10\right\},\left\{\tfrac{8}{5},\tfrac{12}{5},\tfrac{16}{5},14\right\},\left\{\tfrac{8}{5},\tfrac{16}{5},12,14\right\},\left\{\tfrac{8}{5},\tfrac{14}{5},\tfrac{16}{5},14\right\},\left\{\tfrac{8}{5},\tfrac{16}{5},6,14\right\},\left\{\tfrac{16}{5},\tfrac{24}{5},6,14\right\},\left\{\tfrac{12}{5},\tfrac{16}{5},8,14\right\}$
$\left\{\tfrac{16}{5},8,12,14\right\},\left\{\tfrac{14}{5},\tfrac{16}{5},8,14\right\},\left\{\tfrac{16}{5},6,8,14\right\},\left\{\tfrac{12}{5},\tfrac{16}{5},\tfrac{18}{5},14\right\},\left\{\tfrac{12}{5},\tfrac{16}{5},10,14\right\},\left\{\tfrac{16}{5},\tfrac{18}{5},12,14\right\},\left\{\tfrac{16}{5},10,12,14\right\},\left\{\tfrac{14}{5},\tfrac{16}{5},\tfrac{18}{5},14\right\}$
$\left\{\tfrac{14}{5},\tfrac{16}{5},10,14\right\},\left\{\tfrac{16}{5},\tfrac{18}{5},6,14\right\},\left\{\tfrac{16}{5},6,10,14\right\},\left\{\tfrac{8}{5},\tfrac{12}{5},\tfrac{16}{5},4\right\},\left\{\tfrac{8}{5},\tfrac{16}{5},4,12\right\},\left\{\tfrac{8}{5},\tfrac{14}{5},\tfrac{16}{5},4\right\},\left\{\tfrac{8}{5},\tfrac{16}{5},4,6\right\}$
$\left\{\tfrac{12}{5},\tfrac{16}{5},4,8\right\},\left\{\tfrac{16}{5},4,8,12\right\},\left\{\tfrac{14}{5},\tfrac{16}{5},4,8\right\},\left\{\tfrac{16}{5},4,6,8\right\},\left\{\tfrac{12}{5},\tfrac{16}{5},\tfrac{18}{5},4\right\},\left\{\tfrac{12}{5},\tfrac{16}{5},4,10\right\},\left\{\tfrac{16}{5},\tfrac{18}{5},4,12\right\}$
$\left\{\tfrac{16}{5},4,10,12\right\},\left\{\tfrac{14}{5},\tfrac{16}{5},\tfrac{18}{5},4\right\},\left\{\tfrac{14}{5},\tfrac{16}{5},4,10\right\},\left\{\tfrac{16}{5},\tfrac{18}{5},4,6\right\},\left\{\tfrac{16}{5},4,6,10\right\}$
$\tfrac{24}{7}$$\left\{\tfrac{8}{7},\tfrac{12}{7},\tfrac{18}{7},\tfrac{24}{7}\right\},\left\{\tfrac{8}{7},\tfrac{12}{7},\tfrac{24}{7},6\right\},\left\{\tfrac{8}{7},\tfrac{18}{7},\tfrac{24}{7},12\right\},\left\{\tfrac{8}{7},\tfrac{24}{7},6,12\right\},\left\{\tfrac{8}{7},\tfrac{18}{7},\tfrac{24}{7},14\right\},\left\{\tfrac{8}{7},\tfrac{24}{7},6,14\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{8}{7},\tfrac{20}{7},\tfrac{24}{7},14\right\}}}$
$\left\{\tfrac{12}{7},\tfrac{18}{7},\tfrac{24}{7},8\right\},\left\{\tfrac{12}{7},\tfrac{24}{7},6,8\right\},\left\{\tfrac{18}{7},\tfrac{24}{7},8,12\right\},\left\{\tfrac{24}{7},6,8,12\right\},\left\{\tfrac{18}{7},\tfrac{24}{7},8,14\right\},\left\{\tfrac{24}{7},6,8,14\right\},\left\{\tfrac{12}{7},\tfrac{18}{7},\tfrac{24}{7},18\right\}$
$\left\{\tfrac{12}{7},\tfrac{24}{7},6,18\right\},\left\{\tfrac{18}{7},\tfrac{24}{7},12,18\right\},\left\{\tfrac{24}{7},6,12,18\right\},\left\{\tfrac{18}{7},\tfrac{24}{7},\tfrac{30}{7},14\right\},\left\{\tfrac{18}{7},\tfrac{24}{7},14,18\right\},\left\{\tfrac{24}{7},6,14,18\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{20}{7},\tfrac{24}{7},14,18\right\}}}$
Table 10: The genuinely rank-$4$ sets of scaling dimensions with $\tfrac{30}{13}\leq\Delta_{\mathrm{new}}\leq\tfrac{24}{7}$.
$\Delta_{\mathrm{new}}$$\{\Delta_{1},\Delta_{2},\Delta_{3},\Delta_{4}\}$
$\tfrac{15}{4}$$\left\{\tfrac{5}{4},\tfrac{3}{2},\tfrac{15}{4},14\right\},\left\{\tfrac{5}{4},\tfrac{3}{2},3,\tfrac{15}{4}\right\},\left\{\tfrac{5}{4},\tfrac{3}{2},\tfrac{15}{4},18\right\},\left\{\tfrac{5}{4},\tfrac{15}{4},9,14\right\},\left\{\tfrac{5}{4},3,\tfrac{15}{4},9\right\},\left\{\tfrac{5}{4},\tfrac{15}{4},9,18\right\},\left\{\tfrac{5}{4},\tfrac{15}{4},8,14\right\},\left\{\tfrac{5}{4},3,\tfrac{15}{4},8\right\}$
$\left\{\tfrac{5}{4},\tfrac{15}{4},8,18\right\},\left\{\tfrac{3}{2},\tfrac{15}{4},5,14\right\},\left\{\tfrac{3}{2},3,\tfrac{15}{4},5\right\},\left\{\tfrac{3}{2},\tfrac{15}{4},5,18\right\},\left\{\tfrac{15}{4},5,9,14\right\},\left\{3,\tfrac{15}{4},5,9\right\},\left\{\tfrac{15}{4},5,9,18\right\},\left\{\tfrac{15}{4},5,8,14\right\}$
$\left\{3,\tfrac{15}{4},5,8\right\},\left\{\tfrac{15}{4},5,8,18\right\},\left\{\tfrac{3}{2},\tfrac{15}{4},\tfrac{9}{2},14\right\},\left\{\tfrac{3}{2},\tfrac{15}{4},12,14\right\},\left\{\tfrac{3}{2},3,\tfrac{15}{4},\tfrac{9}{2}\right\},\left\{\tfrac{3}{2},3,\tfrac{15}{4},12\right\},\left\{\tfrac{3}{2},\tfrac{15}{4},\tfrac{9}{2},18\right\},\left\{\tfrac{3}{2},\tfrac{15}{4},12,18\right\}$
$\left\{\tfrac{15}{4},\tfrac{9}{2},9,14\right\},\left\{\tfrac{15}{4},9,12,14\right\},\left\{3,\tfrac{15}{4},\tfrac{9}{2},9\right\},\left\{3,\tfrac{15}{4},9,12\right\},\left\{\tfrac{15}{4},\tfrac{9}{2},9,18\right\},\left\{\tfrac{15}{4},9,12,18\right\}$
$\left\{\tfrac{15}{4},\tfrac{9}{2},8,14\right\},\left\{\tfrac{15}{4},8,12,14\right\},\left\{3,\tfrac{15}{4},\tfrac{9}{2},8\right\},\left\{3,\tfrac{15}{4},8,12\right\},\left\{\tfrac{15}{4},\tfrac{9}{2},8,18\right\},\left\{\tfrac{15}{4},8,12,18\right\}$
$\tfrac{30}{7}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{8}{7},\tfrac{18}{7},\tfrac{20}{7},\tfrac{30}{7}\right\}}},\left\{\tfrac{8}{7},\tfrac{18}{7},\tfrac{30}{7},12\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{8}{7},\tfrac{20}{7},\tfrac{30}{7},8\right\}}},\left\{\tfrac{8}{7},\tfrac{30}{7},8,12\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{18}{7},\tfrac{24}{7},\tfrac{30}{7},14\right\}}},\left\{\tfrac{18}{7},\tfrac{30}{7},12,14\right\}$
$\left\{\tfrac{30}{7},8,12,14\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{18}{7},\tfrac{20}{7},\tfrac{30}{7},18\right\}}},\left\{\tfrac{18}{7},\tfrac{30}{7},12,18\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{20}{7},\tfrac{30}{7},8,18\right\}}},\left\{\tfrac{30}{7},8,12,18\right\}$
$\tfrac{24}{5}$${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{6}{5},\tfrac{12}{5},\tfrac{16}{5},\tfrac{24}{5}\right\}}},\left\{\tfrac{6}{5},\tfrac{12}{5},\tfrac{24}{5},8\right\},\left\{\tfrac{6}{5},\tfrac{12}{5},\tfrac{18}{5},\tfrac{24}{5}\right\},\left\{\tfrac{6}{5},\tfrac{12}{5},\tfrac{24}{5},18\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{6}{5},\tfrac{16}{5},\tfrac{24}{5},12\right\}}},\left\{\tfrac{6}{5},\tfrac{24}{5},8,12\right\},\left\{\tfrac{6}{5},\tfrac{18}{5},\tfrac{24}{5},12\right\}$
$\left\{\tfrac{6}{5},\tfrac{24}{5},12,18\right\},\left\{\tfrac{6}{5},\tfrac{24}{5},8,14\right\},\left\{\tfrac{6}{5},\tfrac{18}{5},\tfrac{24}{5},14\right\},\left\{\tfrac{6}{5},\tfrac{24}{5},14,18\right\},\left\{\tfrac{12}{5},\tfrac{24}{5},6,8\right\},\left\{\tfrac{12}{5},\tfrac{18}{5},\tfrac{24}{5},6\right\},\left\{\tfrac{12}{5},\tfrac{24}{5},6,18\right\}$
$\left\{\tfrac{24}{5},6,8,12\right\},\left\{\tfrac{18}{5},\tfrac{24}{5},6,12\right\},\left\{\tfrac{24}{5},6,12,18\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{16}{5},\tfrac{24}{5},6,14\right\}}},\left\{\tfrac{24}{5},6,8,14\right\},\left\{\tfrac{18}{5},\tfrac{24}{5},6,14\right\},\left\{\tfrac{24}{5},6,14,18\right\}$
$\tfrac{16}{3}$$\left\{\tfrac{4}{3},\tfrac{8}{3},\tfrac{10}{3},\tfrac{16}{3}\right\},\left\{\tfrac{4}{3},\tfrac{8}{3},\tfrac{16}{3},14\right\},\left\{\tfrac{4}{3},\tfrac{8}{3},4,\tfrac{16}{3}\right\},\left\{\tfrac{4}{3},\tfrac{10}{3},\tfrac{16}{3},8\right\},\left\{\tfrac{4}{3},\tfrac{16}{3},8,14\right\},\left\{\tfrac{4}{3},4,\tfrac{16}{3},8\right\},\left\{\tfrac{4}{3},\tfrac{10}{3},\tfrac{14}{3},\tfrac{16}{3}\right\},\left\{\tfrac{4}{3},\tfrac{10}{3},\tfrac{16}{3},10\right\}$
$\left\{\tfrac{4}{3},\tfrac{14}{3},\tfrac{16}{3},14\right\},\left\{\tfrac{4}{3},\tfrac{16}{3},10,14\right\},\left\{\tfrac{4}{3},4,\tfrac{14}{3},\tfrac{16}{3}\right\},\left\{\tfrac{4}{3},4,\tfrac{16}{3},10\right\},\left\{4,\tfrac{16}{3},\tfrac{20}{3},8\right\},\left\{\tfrac{14}{3},\tfrac{16}{3},\tfrac{20}{3},14\right\},\left\{\tfrac{8}{3},\tfrac{10}{3},\tfrac{16}{3},12\right\},\left\{\tfrac{8}{3},\tfrac{16}{3},12,14\right\}$
$\left\{\tfrac{8}{3},4,\tfrac{16}{3},12\right\},\left\{\tfrac{10}{3},\tfrac{16}{3},8,12\right\},\left\{\tfrac{16}{3},8,12,14\right\},\left\{4,\tfrac{16}{3},8,12\right\},\left\{\tfrac{10}{3},\tfrac{14}{3},\tfrac{16}{3},12\right\},\left\{\tfrac{10}{3},\tfrac{16}{3},10,12\right\},\left\{\tfrac{14}{3},\tfrac{16}{3},12,14\right\},\left\{\tfrac{16}{3},10,12,14\right\}$
$\left\{4,\tfrac{14}{3},\tfrac{16}{3},12\right\},\left\{4,\tfrac{16}{3},10,12\right\},\left\{\tfrac{8}{3},\tfrac{10}{3},\tfrac{16}{3},6\right\},\left\{\tfrac{8}{3},\tfrac{16}{3},6,14\right\},\left\{\tfrac{8}{3},4,\tfrac{16}{3},6\right\},\left\{\tfrac{10}{3},\tfrac{16}{3},6,8\right\},\left\{\tfrac{16}{3},6,8,14\right\}$
$\left\{4,\tfrac{16}{3},6,8\right\},\left\{\tfrac{10}{3},\tfrac{14}{3},\tfrac{16}{3},6\right\},\left\{\tfrac{10}{3},\tfrac{16}{3},6,10\right\},\left\{\tfrac{14}{3},\tfrac{16}{3},6,14\right\},\left\{\tfrac{16}{3},6,10,14\right\},\left\{4,\tfrac{14}{3},\tfrac{16}{3},6\right\},\left\{4,\tfrac{16}{3},6,10\right\}$
Table 11: The genuinely rank-$4$ sets of scaling dimensions with $\tfrac{15}{4}\leq\Delta_{\mathrm{new}}\leq\tfrac{16}{3}$.
$\Delta_{\mathrm{new}}$$\{\Delta_{1},\Delta_{2},\Delta_{3},\Delta_{4}\}$
$\tfrac{20}{3}$$\left\{\tfrac{4}{3},\tfrac{10}{3},4,\tfrac{20}{3}\right\},\left\{\tfrac{4}{3},\tfrac{10}{3},\tfrac{14}{3},\tfrac{20}{3}\right\},\left\{\tfrac{4}{3},\tfrac{10}{3},\tfrac{20}{3},18\right\},\left\{\tfrac{4}{3},4,\tfrac{20}{3},10\right\},\left\{\tfrac{4}{3},\tfrac{14}{3},\tfrac{20}{3},10\right\},\left\{\tfrac{4}{3},\tfrac{20}{3},10,18\right\},\left\{\tfrac{4}{3},4,\tfrac{20}{3},12\right\}$
$\left\{\tfrac{4}{3},\tfrac{14}{3},\tfrac{20}{3},12\right\},\left\{\tfrac{4}{3},\tfrac{20}{3},12,18\right\},\left\{\tfrac{10}{3},4,\tfrac{20}{3},8\right\},\left\{\tfrac{10}{3},\tfrac{14}{3},\tfrac{20}{3},8\right\},\left\{\tfrac{10}{3},\tfrac{20}{3},8,18\right\},\left\{4,\tfrac{20}{3},8,10\right\},\left\{\tfrac{14}{3},\tfrac{20}{3},8,10\right\}$
$\left\{\tfrac{20}{3},8,10,18\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{4,\tfrac{16}{3},\tfrac{20}{3},8\right\}}},\left\{4,\tfrac{20}{3},8,12\right\},\left\{\tfrac{14}{3},\tfrac{20}{3},8,12\right\},\left\{\tfrac{20}{3},8,12,18\right\},\left\{\tfrac{10}{3},4,\tfrac{20}{3},14\right\},\left\{\tfrac{10}{3},\tfrac{14}{3},\tfrac{20}{3},14\right\},\left\{\tfrac{10}{3},\tfrac{20}{3},14,18\right\}$
$\left\{4,\tfrac{20}{3},10,14\right\},\left\{\tfrac{14}{3},\tfrac{20}{3},10,14\right\},\left\{\tfrac{20}{3},10,14,18\right\},\left\{4,\tfrac{20}{3},12,14\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{\tfrac{14}{3},\tfrac{16}{3},\tfrac{20}{3},14\right\}}},\left\{\tfrac{14}{3},\tfrac{20}{3},12,14\right\},\left\{\tfrac{20}{3},12,14,18\right\}$
$\tfrac{15}{2}$$\left\{\tfrac{3}{2},3,\tfrac{9}{2},\tfrac{15}{2}\right\},\left\{\tfrac{3}{2},3,\tfrac{15}{2},12\right\},\left\{\tfrac{3}{2},3,5,\tfrac{15}{2}\right\},\left\{\tfrac{3}{2},\tfrac{9}{2},\tfrac{15}{2},18\right\},\left\{\tfrac{3}{2},\tfrac{15}{2},12,18\right\},\left\{\tfrac{3}{2},5,\tfrac{15}{2},18\right\},\left\{\tfrac{3}{2},\tfrac{9}{2},\tfrac{15}{2},14\right\}$
$\left\{\tfrac{3}{2},\tfrac{15}{2},12,14\right\},\left\{\tfrac{3}{2},5,\tfrac{15}{2},14\right\},\left\{3,\tfrac{9}{2},\tfrac{15}{2},9\right\},\left\{3,\tfrac{15}{2},9,12\right\},\left\{3,5,\tfrac{15}{2},9\right\},\left\{\tfrac{9}{2},\tfrac{15}{2},9,18\right\},\left\{\tfrac{15}{2},9,12,18\right\}$
$\left\{5,\tfrac{15}{2},9,18\right\},\left\{\tfrac{9}{2},\tfrac{15}{2},9,14\right\},\left\{\tfrac{15}{2},9,12,14\right\},\left\{5,\tfrac{15}{2},9,14\right\},\left\{3,\tfrac{9}{2},\tfrac{15}{2},8\right\},\left\{3,\tfrac{15}{2},8,12\right\},\left\{3,5,\tfrac{15}{2},8\right\}$
$\left\{\tfrac{9}{2},\tfrac{15}{2},8,18\right\},\left\{\tfrac{15}{2},8,12,18\right\},\left\{5,\tfrac{15}{2},8,18\right\},\left\{\tfrac{9}{2},\tfrac{15}{2},8,14\right\},\left\{\tfrac{15}{2},8,12,14\right\},\left\{5,\tfrac{15}{2},8,14\right\}$
$15$$\left\{3,5,8,15\right\},\left\{3,5,9,15\right\},\left\{3,8,12,15\right\},\left\{3,9,12,15\right\},\left\{5,8,15,18\right\},\left\{5,9,15,18\right\}$
$\left\{8,12,15,18\right\},\left\{9,12,15,18\right\},\left\{5,8,14,15\right\},\left\{5,9,14,15\right\},\left\{8,12,14,15\right\},\left\{9,12,14,15\right\}$
$16$$\left\{4,6,8,16\right\},\left\{4,6,10,16\right\},\left\{4,8,12,16\right\},\left\{4,10,12,16\right\},\left\{6,8,14,16\right\},\left\{6,10,14,16\right\},\left\{8,12,14,16\right\},\left\{10,12,14,16\right\}$ }
$20$$\left\{4,8,10,20\right\},\left\{4,8,12,20\right\},\left\{4,10,14,20\right\},\left\{4,12,14,20\right\},\left\{8,12,20,24\right\}$
$\left\{8,10,18,20\right\},\left\{8,18,20,30\right\},\left\{8,12,18,20\right\},\left\{10,14,18,20\right\},\left\{12,14,18,20\right\}$
$24$$\left\{6,8,12,24\right\},\left\{6,8,14,24\right\},\left\{6,12,18,24\right\},\left\{6,14,18,24\right\},\left\{12,18,24,30\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{8,12,20,24\right\}}}$
$30$$\left\{8,12,14,30\right\},\left\{8,12,18,30\right\},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{8,18,20,30\right\}}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\left\{12,18,24,30\right\}}}$
Table 12: The genuinely rank-$4$ sets of scaling dimensions with $\tfrac{20}{3}\leq\Delta_{\mathrm{new}}\leq 30$.